### v2

# ProG: A Graph Prompt Learning Benchmark

Abstract
Artificial general intelligence on graphs has shown significant advancements across various applications, yet the traditional ‘Pre-train & Fine-tune’ paradigm faces inefficiencies and negative transfer issues, particularly in complex and few-shot settings. Graph prompt learning emerges as a promising alternative, leveraging lightweight prompts to manipulate data and fill the task gap by reformulating downstream tasks to the pretext. However, several critical challenges still remain: how to unify diverse graph prompt models, how to evaluate the quality of graph prompts, and to improve their usability for practical comparisons and selection. In response to these challenges, we introduce the first comprehensive benchmark for graph prompt learning. Our benchmark integrates SIX pre-training methods and FIVE state-of-the-art graph prompt techniques, evaluated across FIFTEEN diverse datasets to assess performance, flexibility, and efficiency. We also present ‘ProG’, an easy-to-use open-source library that streamlines the execution of various graph prompt models, facilitating objective evaluations. Additionally, we propose a unified framework that categorizes existing graph prompt methods into two main approaches: prompts as graphs and prompts as tokens. This framework enhances the applicability and comparison of graph prompt techniques. The code is available at: https://github.com/sheldonresearch/ProG.
1 Introduction
Recently, artificial general intelligence (AGI) on graphs has emerged as a new trend in various applications like drug design [53, 33], protein prediction [9], social analysis [44, 58], etc. To achieve this vision, one key question is how to learn useful knowledge from non-linear data (like graphs) and how to apply it to various downstream tasks or domains. Classical approaches mostly leverage the ‘Pre-train & Fine-tune’ paradigm, which first designs some pretext with easily access data as the pre-training task for the graph neural network model, and then adjusts partial or entire model parameters to fit new downstream tasks. Although much progress has been achieved, they are still not effective and efficient enough. For example, adjusting the pre-trained model will become very time-consuming with the increase in model complexity [24, 23, 43, 22]. In addition, a natural gap between these pretexts and downstream tasks makes the task transferring very hard, and sometimes may even cause negative transfer [51]. These problems are particularly serious in few-shot settings.
To further alleviate the problems above, Graph Prompt Learning [45] has attracted more attention recently. As shown in Figure 1, graph prompts seek to manipulate downstream data by inserting an additional small prompt graph and then reformulating the downstream task to the pre-training task without changing the pre-trained Graph Neural Network (GNN) model. Since a graph prompt is usually lightweight, tuning this prompt is more efficient than the large backbone model. Compared with graph ‘pre-training & fine-tuning’, which can be seen as a model-level retraining skill, graph ‘pre-training & prompting’ is more customized for data-level operations, which means its potential applications might go beyond task transferring in the graph intelligence area like data quality evaluation, multi-domain alignment, learnable data augmentation, etc. Recently, many graph prompt models have been proposed and presented significant performance in graph learning areas [42, 41, 28, 7, 60, 4], indicating their huge potential towards more general graph intelligence. Despite the enthusiasm around graph prompts in research communities, three significant challenges impede further exploration:
Figure 1: An overview of our benchmark.
  • How can we unify diverse graph prompt models given their varying methodologies? The fragmented landscape of graph prompts hinders systematic research advancement. A unified framework is essential for integrating these diverse methods, creating a cohesive taxonomy to better understand current research and support future studies.
  • How do we evaluate the quality of graph prompts? Assessing the efficiency, power, and flexibility of graph prompts is crucial for understanding their impact and limitations. Currently, there is no standardized benchmark for fair and comprehensive comparison, as existing works have inconsistent experimental setups, varying tasks, metrics, pre-training strategies, and data-splitting techniques, which obstruct a comprehensive understanding of the current state of research.
  • How can we make graph prompt approaches more user-friendly for practical comparison and selection? Despite numerous proposed methods, the lack of an easy-to-use toolkit for creating graph prompts limits their potential applications. The implementation details of existing works vary significantly in programming frameworks, tricks, and running requirements. Developing a standardized library for various graph prompt learning approaches is urgently needed to facilitate broader exploration and application.
In this work, we wish to push forward graph prompt research to be more standardized, provide suggestions for practical choices, and uncover both their advantages and limitations with a comprehensive evaluation. To the best of our knowledge, this is the first benchmark for graph prompt learning. To solve the first challenge, our benchmark treats existing graph prompts as three basic components: prompt tokens, token structure, and insert patterns. According to their detailed implementation, we categorize the current graph prompts into two branches: graph prompt as an additional graph, and graph prompt as tokens. This taxonomy makes our benchmark approach compatible with nearly all existing graph prompting approaches. To solve the second challenge, our benchmark encompasses 6 different classical pre-training methods covering node-level, edge-level, and graph-level tasks, and 5 most representative and state-of-the-art graph prompt methods. Additionally, we include 15 diverse datasets in our benchmark with different scales, covering both homophilic and heterophilic graphs from various domains. We systematically investigate the effectiveness, flexibility, and efficiency of graph prompts under few-shot settings. Compared with traditional supervised methods and ‘pre-training & fine-tuning’ methods, graph prompting approaches present significant improvements. We even observe some negative transfer cases caused by ‘pre-training & fine-tuning’ methods but effectively reversed by graph prompting methods. To solve the third challenge, we develop a unified library for creating various graph prompts. In this way, users can evaluate their models or datasets with less effort. The contributions are summarized as follows:
  • We propose the first comprehensive benchmark for graph prompt learning. The benchmark integrates 6 most used pre-training methods and 5 state-of-the-art graph prompting methods with 15 diverse graph datasets. We systematically evaluate the effectiveness, flexibility, and efficiency of current graph prompt models to assist future research efforts. (Detailed in Section 5)
  • We offer ‘ProG’ [^1] Our code is available at https://github.com/sheldonresearch/ProG, an easy-to-use open-source library to conduct various graph prompt models (Section 3). We can get rid of the distraction from different implementation skills and reveal the potential benefits/shortages of these graph prompts in a more objective and fair setting.
  • We propose a unified view for current graph prompt methods (Section 2). We decompose graph prompts into three components: prompt tokens, token structure, and insert patterns. With this framework, we can easily group most of the existing work into two categories (prompt as graph, and prompt as tokens). This view serves as the infrastructure of our developed library and makes our benchmark applicable to most of the existing graph prompt works.
2 Preliminaries
Understand Graph Prompt Nature in A Unified View. Graph prompt learning aims to learn suitable transformation operations for graphs or graph representations to reformulate the downstream task to the pre-training task. Let $t(\cdot)$ be any graph-level transformation (e.g., "modifying node features", "augmenting original graphs", etc.), and $\psi^{*}(\cdot)$ be a frozen pre-trained graph model. For any graph $\mathcal{G}$ with adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$ and node feature matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ where $N$ denotes the node number and $d$ is feature number. It has been proven [42, 7] that we can always learn an appropriate prompt module $\mathcal{P}$, making them can be formulated as the following formula:
$\psi^{*}(\mathcal{P}(\mathbf{X},\mathbf{A},\mathbf{P},\mathbf{A}_{inner},\mathbf{A}_{cross}))=\psi^{*}(t(\mathbf{X},\mathbf{A}))+O_{\mathcal{P}\psi}$ (1)
Here $\mathbf{P}\in\mathbb{R}^{K\times d}$ is the learnable representations of $K$ prompt tokens, $\mathbf{A}_{inner}\in\{0,1\}^{K\times K}$ indicate token structures and $\mathbf{A}_{cross}\in\{0,1\}^{K\times N}$ is the inserting patterns, which indicates how each prompt token connects to each node in the original graph. Prompt module $\mathcal{P}$ takes the input graph and then outputs an integrated graph with a graph prompt. This equation indicates that we can learn an appropriate prompt module $\mathcal{P}$ applied to the original graph to imitate any graph manipulation. $O_{\mathcal{P}\psi}$ denotes the error bound between the manipulated graph and the prompting graph w.r.t. their representations from the pre-trained graph model, which can be seen as a measurement of graph prompt flexibility.
Task Definition (Few-shot Graph Prompt Learning). Given a graph $\mathcal{G}$, a frozen pre-trained graph model $\psi^{*}$, and labels $\mathcal{Y}$, the loss function is to optimize the task loss $\mathcal{L}_{Task}$ as follows:
$\mathcal{L}_{\mathcal{P}}=\mathcal{L}_{Task}(\psi^{*}(\mathcal{P}(\mathbf{X},\mathbf{A},\mathbf{P},\mathbf{A}_{inner},\mathbf{A}_{cross})),\mathcal{Y})$
During the few-shot training phase, the labeled samples are very scarce. Specifically, given a set of classes $C$, each class $C_{i}$ has only $k$ labeled examples. Here, $k$ is a small number (e.g., $k\leq 10$), and the total number of labeled samples is given by $k\times|C|=|\mathcal{Y}|$, where $|C|$ is the number of classes. This setup is commonly referred to as $k$-shot classification [28]. In this paper, our evaluation includes both node and graph classification tasks.
Evaluation Questions. We are particular interested in the effectiveness, flexibility, and efficiency of graph prompt learning methods. Specifically, in effectiveness, we wish to figure out how effective are different graph prompt methods on various tasks (Section 5.1); and how well graph prompting methods overcome the negative transferring caused by pre-training strategies (Section 5.2). In flexibility, we wish to uncover how powerful different graph prompt methods are to simulate data operations (Section 5.3). In efficiency, we wish to know how efficient are these graph prompt methods in terms of time and space cost (Section 5.4).
3 Benchmark Methods
In general, graph prompt learning aims to transfer pre-trained knowledge with the expectation of achieving better positive transfer effects in downstream tasks compared to classical ‘Pre-train & Fine-tune’ methods. To conduct a comprehensive comparison, we first consider using a supervised method as the baseline for judging positive transfer (For example, if a ‘pre-training and fine-tuning’ approach can not beat supervised results, one negative transfer case is observed). We introduce various pre-training methods to generate pre-trained knowledge, upon which we perform fine-tuning as the comparison for graph prompts. Finally, we apply current popular graph prompt learning methods with various pre-training methods to investigate their knowledge transferability. The details of these baselines are as follows (Further introduction can be referred to Appendix A):
  • Supervised Method: In this paper, we utilize Graph Convolutional Network (GCN) [20] as the baseline, which is one of the classical and effective graph models based on graph convolutional operations [45, 61] and also the backbone for both ‘Pre-train & Fine-tune’ and graph prompt learning methods. We also explore more kinds of graph models like GraphSAGE [12], GAT [48], and Graph Transformer [38], and put the results in Appendix E.
  • ‘Pre-train & Fine-tune’ Methods: We select 6 mostly used pre-training methods covering node-level, edge-level, and graph-level strategies. For node-level, we consider DGI [49] and GraphMAE [13], where DGI maximizes the mutual information between node and graph representations for informative embeddings and GraphMAE learns deep node representations by reconstructing masked features. For edge-level, we introduce EdgePreGPPT [41] and EdgePreGprompt [28], where EdgePreGPPT calculates the dot product as the link probability of node pairs and EdgePreGprompt samples triplets from label-free graphs to increase the similarity between the contextual subgraphs of linked pairs while decreasing the similarity of unlinked pairs. For graph-level, we involve GCL [59], SimGRACE [54], where GCL maximizes agreement between different graph augmentations to leverage structural information and SimGRACE tries to perturb the graph model parameter spaces and narrow down the gap between different perturbations for the same graph.
  • Graph Prompt Learning Methods: With our proposed unified view, current popular graph prompt learning methods can be easily classified into two types, ‘Prompt as graph’ and ‘Prompt as token’ types. Specifically, the ‘Prompt as graph’ type means a graph prompt has multiple prompt tokens with inner structure and insert patterns, and we introduce All-in-one [42], which aims to learn a set of insert patterns for original graphs via learnable graph prompt modules. For the ‘Prompt as token’ type, we consider involving GPPT [41], Gprompt [28], GPF and GPF-plus [7]. Concretely, GPPT defines graph prompts as additional tokens that contain task tokens and structure tokens that align downstream node tasks and link prediction pretext. Gprompt focuses on inserting the prompt vector into the graph pooling by element-wise multiplication. GPF and GPF-plus mainly add soft prompts to all node features of the input graph.
Figure 2: An overview of ProG
ProG: A Unified Library for Graph Prompts. We offer our developed graph prompt library ProG as shown in Figure 2. In detail, the library first offers a Model Backbone module, including many widely-used graph models, which can be used for supervised learning independently. Then, it designs a Pre-training module involving many pre-training methods from various level types, which can pre-train the graph model initialized from the model backbone to preserve pre-trained knowledge and further fine-tune the pre-trained graph model. Finally, it seamlessly incorporates different graph prompt learning methods together into a unified Prompting module, which can freeze and execute prompt tuning for the pre-trained graph model from the pre-training module. For a fair comparison, it defines an Evaluation module, including comprehensive metrics, batch evaluator, and dynamic dispatcher, to ensure the same settings across different selected methods. Many essential operations are fused into the Utils module for repeated use, and all the data-related operations are included in the Data module for user-friendly operations to use the library. Additionally, the Tutorial module introduces the key functions that ProG offers.
4 Benchmark Settings
4.1 Datasets
Table 1: Statistics of all datasets. N indicates node classification, and G indicates graph classification.
Dataset Graphs Avg.nodes Avg.edges Features Node classes Task (N / G) Category
Cora 1 2,708 5,429 1,433 7 N Homophilic
Pubmed 1 19,717 88,648 500 3 N Homophilic
CiteSeer 1 3,327 9,104 3,703 6 N Homophilic
Actor 1 7600 30019 932 5 N Heterophilic
Wisconsin 1 251 515 1703 5 N Heterophilic
Texas 1 183 325 1703 5 N Heterophilic
ogbn-arxiv 1 169,343 1,166,243 128 40 N Homophilic & Large scale
Dataset Graphs Avg.nodes Avg.edges Features Graph classes Task (N / G) Domain
MUTAG 188 17.9 19.8 7 2 G small molecule
IMDB-BINARY 1000 19.8 96.53 0 2 G social network
COLLAB 5000 74.5 2457.8 0 3 G social network
PROTEINS 1,113 39.1 72.8 3 2 G proteins
ENZYMES 600 32.6 62.1 3 6 G proteins
DD 1,178 284.1 715.7 89 2 G proteins
COX2 467 41.2 43.5 3 2 G small molecule
BZR 405 35.8 38.4 3 2 G small molecule
To figure out how adaptive existing graph prompts on different graphs, we conduct our experiments across 15 datasets on node-level or graph-level tasks, providing a broad and rigorous testing ground for algorithmic comparisons. As shown in Table 4.1, we include 7 node classification datasets, covering homophilic datasets (Cora, Citeseer, PubMed, ogbn-arxiv) [37, 30, 14], heterophilic datasets (Wisconsin, Texas, Actor) [32, 47], and large-scale datasets (ogbn-arxiv). Additionally, we consider 8 graph classification datasets from different domains, including social networks (IMDB-B, COLLAB) [57], biological datasets (ENZYMES, PROTEINS, DD) [6, 2, 52], and small molecule datasets (MUTAG, COX2, BZR) [21, 36]. More details on these datasets can be found in Appendix C.
4.2 Implementation Details
Data Split. For the node task, we use 90% data as the test set. For the graph task, we use 80% data as the test set. To ensure the robustness of our findings, we repeat the sampling process five times to construct k-shot tasks for both node and graph tasks. We then report the average and standard deviation over these five results.
Metrics. Following existing multi-class classification benchmarks, we select Accuracy, Macro F1 Score, and AUROC (Area Under the Receiver Operating Characteristic Curve) as our main performance metrics. The Accuracy results are used to provide a straightforward indication of the model’s overall performance in Section 5. F1 Score and AUROC results are provided in Appendix D.
Prompt Adaptation to Specific Tasks. We observe that certain prompt methods, such as GPPT, are only applicable to node classification, while GPF and GPF-plus are only suitable for graph classification. To this end, we design graph task tokens to replace the original node task tokens to make GPPT also applicable in graph-level tasks. Inspired by All-in-one and Gprompt, we utilize induced graphs in our benchmark to adapt GPF and GPF-plus for the node task. The concrete adaptation designs are shown in Appendix A.
Hyperparameter Optimization. To manage the influence of hyperparameter selection and maintain fairness, the evaluation process is standardized both with and without hyperparameter adjustments. Initially, default hyperparameters as specified in the original publications are used. To ensure fairness during hyperparameter tuning, random search is then employed to optimize the settings. In each trial across different datasets, a set of hyperparameters is randomly selected from a predefined search space for each model. For additional details on metrics, default hyperparameters, search spaces, and other implementation aspects, please see Appendix D.
5 Experimental Results
5.1 Overall Performance of Graph Prompt Learning
In Table 2 and Table 3, we present the optimal experimental results of the supervised methods, classical ‘Pre-train & Fine-tune’ methods, and various graph prompt methods on 1-shot node/graph classification tasks across 15 datasets, where the evaluation metric is accuracy (Complete results can be referred to Appendix E). Here the ’optimal’ results mean that the best result is shown from six classical ‘Pre-train & Fine-tune’ methods or from six pretraining-based variants of each graph prompt method. This presentation style can help reflect the upper bounds of classical ‘Pre-train & Fine-tune’ methods and graph prompt methods, better highlighting the advantages of graph prompts.
From these results, we can observe consistent advantages of various graph prompt learning methods. Almost all graph prompt methods consistently outperform the Supervised method on both node-level and graph-level tasks, demonstrating positive transfer and superior performance. Specifically, graph prompt methods generally surpass classical ‘Pre-train & Fine-tune’ methods across most datasets, showcasing enhanced knowledge transferability. In node-level tasks, GPF-plus, which focuses on modifying node features via learnable prompt features, achieves the best results on 4 out of 7 datasets with significant effects, attributed to its inherent ability to easily adapt to different pre-training methods [7]. In contrast, All-in-one, which concentrates on modifying graphs via learnable subgraphs, achieves the best results on 7 out of 8 datasets on graph-level tasks.
Table 2: Performance on 1-shot node classification. The best results for each dataset are highlighted in bold with a dark red background. The second-best are underlined with a light red background.
[…]
Table 3: Performance on 1-shot graph classification. The best results for each dataset are highlighted in bold with a dark red background. The second-best are underlined with a light red background.
[…]
Table 4: 1-shot node classification performance comparison across different datasets of GPF-plus. ↑/↓ indicates the positive/negative transfer compared to the supervised baseline. ↓(%) means the negative transfer rate, computed by dividing the number of ↓ by the number of datasets.
[…]
Supervised 26.56±5.55 (-) 21.78±7.32 (-) 39.37±16.34 (-) 41.60±3.10 (-) 37.97±5.80 (-) 20.57±4.47 (-) 10.99±3.19 (-) 0.0%
DGI 33.15±7.84 (↑) 21.64±3.92 (↓) 42.01±12.54 (↑) 37.49±7.56 (↓) 45.31±5.01 (↑) 19.76±3.53 (↓) 7.21±2.91 (↓) 57.0%
+ GPF_plus 17.29±6.18 (↓) 26.60±13.24 (↑) 34.02±11.94 (↓) 76.34±6.26 (↑) 79.54±1.89 (↑) 22.42±9.66 (↑) 16.83±10.02 (↑) 29%
GraphMAE 32.93±3.17 (↑) 21.26±3.57 (↓) 42.99±14.25 (↑) 36.80±7.17 (↓) 37.81±8.62 (↓) 19.86±2.70 (↓) 12.35±3.60 (↑) 43.0%
+ GPF_plus 54.26±7.48 (↑) 59.67±11.87 (↑) 46.64±18.57 (↑) 89.17±9.79 (↑) 94.09±4.40 (↑) 26.58±7.84 (↑) 49.81±2.62 (↑) 0.0%
EdgePreGPPT 38.12±5.29 (↑) 18.09±5.39 (↓) 46.74±14.09 (↑) 35.31±9.31 (↓) 47.66±2.37 (↑) 19.17±2.53 (↓) 16.21±3.82 (↑) 43.0%
+ GPF_plus 28.49±18.73 (↓) 28.04±14.31 (↑) 46.51±15.84 (↑) 91.03±4.11 (↑) 92.14±5.46 (↑) 29.32±8.56 (↑) 71.98±12.23 (↑) 14.0%
EdgePreGprompt 35.57±5.83 (↑) 22.28±3.80 (↑) 41.50±7.54 (↑) 40.69±4.13 (↓) 40.62±7.95 (↑) 20.74±4.16 (↑) 14.83±2.38 (↑) 14.0%
+ GPF_plus 55.77±10.30 (↑) 49.43±8.21 (↑) 42.79±18.18 (↑) 85.44±2.23 (↑) 87.70±3.77 (↑) 22.68±3.64 (↑) 57.44±6.95 (↑) 0.0%
GCL 52.61±1.73 (↑) 27.02±4.31 (↑) 42.49±11.29 (↑) 33.94±7.74 (↓) 40.31±13.68 (↑) 20.19±1.98 (↓) 4.65±1.19 (↓) 43.0%
+ GPF_plus 34.18±17.71 (↓) 28.86±22.88 (↑) 37.02±11.29 (↓) 68.57±8.77 (↑) 95.83±4.19 (↑) 22.82±4.99 (↑) 32.11±4.86 (↑) 29.0%
SimGRACE 40.40±4.66 (↑) 35.05±4.37 (↑) 37.59±8.17 (↓) 37.37±3.68 (↓) 46.88±4.64 (↑) 19.78±1.89 (↓) 8.13±3.26 (↓) 57.0%
+ GPF_plus 21.33±14.86 (↓) 24.61±21.21 (↓) 35.90±9.06 (↓) 73.64±5.25 (↑) 90.74±3.12 (↑) 20.51±4.24 (↑) 46.71±3.17 (↑) 43.0%
Table 5: 1-shot graph classification performance comparison across different datasets of All-in-one.
[…]
Supervised 57.30±0.98 (-) 47.23±0.61 (-) 56.36±7.97 (-) 65.20±6.70 (-) 20.58±2.00 (-) 27.08±11.94 (-) 25.80±6.53 (-) 55.33±6.22 (-) 0.0%
DGI 57.32±0.90 (↑) 42.22±0.73 (↓) 64.65±2.10 (↑) 64.13±7.90 (↓) 17.83±1.88 (↓) 29.44±9.68 (↑) 26.48±7.61 (↑) 57.15±4.32 (↑) 38.0%
 + All-in-one 60.07±4.81 (↑) 39.56±5.00 (↓) 62.58±7.07 (↑) 75.73±7.75 (↑) 23.96±1.45 (↑) 50.72±9.93 (↑) 79.20±1.65 (↑) 55.97±6.52 (↑) 13.0%
GraphMAE 57.70±1.13 (↑) 48.10±0.23 (↑) 63.57±3.57 (↑) 65.20±5.00 (-) 22.21±2.79 (↑) 28.47±14.72 (↑) 25.80±6.53 (-) 57.54±4.41 (↑) 0.0%
 + All-in-one 52.62±3.04 (↓) 40.82±14.63 (↓) 66.49±6.26 (↑) 72.27±8.87 (↑) 23.21±1.72 (↑) 56.68±7.38 (↑) 75.37±5.99 (↑) 58.77±1.05 (↑) 25.0%
EdgePreGPPT 57.20±0.85 (↓) 47.14±0.55 (↓) 58.27±10.66 (↑) 64.27±4.73 (↓) 19.79±2.17 (↓) 27.83±13.44 (↑) 72.10±14.30 (↑) 52.82±9.38 (↓) 63.0%
 + All-in-one 59.12±0.77 (↑) 42.74±4.65 (↓) 65.71±5.49 (↑) 79.87±5.34 (↑) 20.92±2.04 (↑) 60.27±16.97 (↑) 64.20±19.74 (↑) 56.24±2.46 (↑) 13.0%
EdgePreGprompt 57.35±0.92 (↑) 47.20±0.53 (↓) 61.84±2.59 (↑) 62.67±2.67 (↓) 19.75±2.33 (↓) 27.13±12.05 (↑) 29.44±11.20 (↑) 56.16±5.10 (↑) 38.0%
 + All-in-one 53.78±2.82 (↓) 42.87±6.19 (↓) 61.82±7.53 (↑) 75.47±5.00 (↑) 21.88±0.56 (↑) 49.06±5.53 (↑) 66.60±17.36 (↑) 57.60±4.37 (↑) 25.0%
GCL 57.75±1.02 (↑) 39.62±0.63 (↓) 63.44±3.64 (↑) 65.07±8.38 (↓) 23.96±1.99 (↑) 53.14±21.32 (↑) 29.07±7.00 (↑) 60.62±1.56 (↑) 25.0%
 + All-in-one 58.75±0.80 (↑) 51.66±2.60 (↑) 64.36±7.30 (↑) 70.53±8.58 (↑) 19.46±2.85 (↓) 52.55±13.51 (↑) 50.93±16.83 (↑) 59.72±1.52 (↑) 13.0%
SimGRACE 57.33±0.96 (↑) 46.89±0.42 (↓) 60.07±3.21 (↑) 65.47±5.89 (↑) 19.71±1.76 (↓) 76.19±5.41 (↑) 28.48±6.49 (↑) 53.23±9.71 (↓) 38.0%
 + All-in-one 58.83±0.85 (↑) 47.60±3.90 (↑) 61.17±1.73 (↑) 65.47±6.91 (↑) 22.50±1.56 (↑) 76.14±5.51 (↑) 69.88±18.06 (↑) 58.26±1.18 (↑) 0.0%
Similar observations can also be found in 3-shot and 5-shot results in the optimal and complete versions, which can be checked in Appendix E. Additionally, for comprehensive, we also offer the curve of the shot number from 1 to 10 and the accuracy in Appendix E.
5.2 Impact of Pre-training Methods
From Table 2, we further find that on the Wisconsin dataset, the classical Pre-train & Fine-tune method consistently exhibited negative transfer, whereas most graph prompt methods effectively transferred upstream pre-training knowledge, achieving significant results and positive transfer. Therefore, it is crucial to explore the effectiveness of various pre-training methods, especially for graph prompt learning. Table 4 and Table 5 show the experimental results of six variants of GPF-plus/All-in-one and the corresponding ‘Pre-train & Fine-tune’ methods. Concretely, we have the following key observations:
Consistency in Pretext and Downstream Tasks. In node classification tasks, the pre-trained knowledge learned by the node-level pre-training method (e.g., GraphMAE) shows superior transferability to downstream tasks using GPF-plus, achieving better results compared to traditional fine-tuning methods. Conversely, in graph classification tasks, the graph-level pre-training method (e.g., SimGRACE) demonstrates better transferability with All-in-one, resulting in more significant outcomes. These observations indicate that although graph prompts may perform better, they are also selective to specific pre-training tasks. In general, node-level pre-training is more suitable for node-level tasks, while graph-level pre-training is more effective for graph-level tasks.
Mitigation of Negative Transfer. Classical pre-trained models often experience significant negative transfer issues during fine-tuning. In contrast, prompt tuning can effectively handle these problems. For example, in node classification, GPF-plus can reduce the negative transfer rate from 43% to 0%. Similarly, in graph classification, All-in-one successfully resolves negative transfer issues, reducing the rate from 38% to 0%. This demonstrates the efficacy of graph prompt methods, such as GPF-plus and All-in-one, in mitigating negative transfer, ensuring more reliable and accurate knowledge transfer in both node-level and graph-level tasks.
(a) GPF-plus (1-shot node classification Task).
(b) All-in-one (1-shot graph classification Task).
Figure 3: Head map of GPF-plus and All-in-one on node-level and graph-level tasks across various pre-training methods and datasets. The green interval shows that the graph prompt learning method performs worse than ‘Pre-train & Fine-tune’. The yellow-to-red interval shows that the graph prompt learning method performs better. An upward arrow indicates that the ‘Pre-train & Fine-tune’ method experiences negative transfer, while the graph prompt learning method can alleviate it significantly.
Figure 3(a) and Figure 3(b) respectively show the heat maps of Table 4 and Table 5, respectively, for help better understanding the impact of pre-training methods (Other heat maps of other graph prompt learning methods can be referred to Appendix E).
5.3 Flexibility Analysis
Table 6: Error bound between different prompt methods, RED(%): average reduction of each method to the original error.
Prompt Model Drop Drop Mask RED (%)
Nodes Edges Feature
ori_error 0.4862 0.0713 0.6186
All-in-one 0.0200 0.0173 0.0207 95.06 $\downarrow$
Gprompt 0.0218 0.0141 0.0243 94.88 $\downarrow$
GPF-plus 0.0748 0.0167 0.1690 77.85 $\downarrow$
GPF 0.0789 0.0146 0.1858 76.25 $\downarrow$
We leverage error bound, defined by Equation 1, to measure the flexibility of graph prompt methods [42]. The error bound quantifies the difference in graph representations between the original graph and the manually manipulated graph restored by graph prompt methods. These manipulations include dropping nodes, dropping edges, and masking features. Table 5.3 presents the specific results. The experimental results indicate that all graph prompt methods significantly reduced the error compared to the original error across all types of manipulations. Notably, All-in-one and Gprompt achieved the lowest error bounds, demonstrating their flexibility in adapting to graph transformations. Specifically, All-in-one achieved a REDof 95.06, indicating a 95.06% reduction in original error, underscoring its comprehensive approach to capturing graph transformations. These findings highlight the importance of flexibility in prompt design for effective knowledge transfer. Flexible prompts that can adapt to various graph transformations enable more powerful graph data operation capability, thereby improving the performance of downstream tasks. Moreover, this flexibility is inspiring for more customized graph data augmentation, indicating more potential for other applications.
5.4 Efficiency Analysis
We evaluate the efficiency and training duration of various graph prompt methods across tasks, focusing on node classification in Cora and Wisconsin and graph classification in BZR and DD. GPF-plus achieves high accuracy with relatively low training times despite its larger parameter size on Cora and Wisconsin. In contrast, All-in-One, with a smaller parameter size, has slightly larger training times than GPF-plus due to similarity calculations. Among all the graph prompt methods, GPPT shows lower performance and the longest training duration because of its k-means clustering update in every epoch. Besides, we compare the tunable parameters of graph prompt methods during the training phase with classical ‘Pre-train & Fine-tune’ methods with various backbones for further analyzing the advantages of graph prompt learning. Consider an input graph has $N$ nodes, $M$ edges, $d$ features, along with a graph model which has $L$ layers with maximum layer dimension $D$. Classical backbones (e.g., GCN) have $O(dD+D^{2}(L-1))$ learnable parameter complexity. Other more complex backbones (e.g., graph transformer) may have larger parameters. In the graph prompt learning framework, tuning the prompt parameters with a frozen pre-trained backbone makes training converge faster. For All-in-One, with $K$ tokens and $m$ edges, learnable parameter complexity is $O(Kd)$. Other graph prompt methods also have similar tunable parameter advantages, with specific sizes depending on their designs (e.g., Gprompt only includes a learnable vector inserted into the graph pooling by element-wise multiplication).
Figure 4: Analysis of accuracy, training time and tunable parameters of various graph prompt methods. Note that the gray area enclosed by the dashed line represents the scale of the tunable parameters.
5.5 Adaptability of Graph Prompts on Different Graphs.
Besides the main observations mentioned before, we here further discuss the adaptability of graph prompts on various kinds of graphs. Specifically, considering graph prompts performance on various domains as shown in Table 3, we can find that the adaptability to various domains of graph prompt methods, particularly those of the Prompt as Token type, is not always good, where graph prompt methods even may encounter negative transfer issues. GPPT, in particular, faces significant challenges, likely because first-order subgraphs fail to comprehensively capture graph patterns, introducing noise and affecting transferability. For Large-scale Datasets like node-level dataset ogbn-arxiv in Table 2, the transferability of graph prompt methods involving subgraph operations, such as GPPT and All-in-one, requires further enhancement. Furthermore, regarding the relatively large graph-level dataset DD in Table 3, all graph prompt methods achieve positive transfer and outperform Pre-train & Fine-tune. Despite this success, the magnitude of improvement indicates that the transferability on graph-level large-scale datasets also requires further enhancement. For Homophilic & Heterophilic Datasets in Table 2, most graph prompt methods can handle both homophilic and heterophilic graph datasets effectively, with the exception of GPPT, possibly due to instability caused by the clustering algorithms of GPPT. For Pure Structural Dataset like IMDB-BINARY and COLLAB in Table 3, the ability to handle pure structural datasets of graph prompt methods needs further refinement. In detail, some ‘Pre-train & Fine-tune’ methods can achieve positive transfer on these datasets, while some variants of the All-in-one method even exhibit negative transfer. Further detailed analysis can be referred to Appendix E.
6 Conclusion and Future Plan
In this paper, we introduce ProG, a comprehensive benchmark for comparing various graph prompt learning methods with supervised and ‘Pre-train & Fine-tune’ methods. Our evaluation includes 6 Pre-training methods and 5 graph prompt learning methods across 15 real-world datasets, demonstrating that graph prompt methods generally outperform classical methods. This performance is attributed to the high flexibility and low complexity of graph prompt methods. Our findings highlight the promising future of graph prompt learning research and its challenges. By making ProG open-source, we aim to promote further research and better evaluations of graph prompt learning methods.
As the empirical analysis in this paper, graph prompt learning shows its significant potential beyond task transferring, though our current evaluation focuses on node/graph-level tasks. We view ProG as a long-term project and are committed to its continuous development. Our future plans include expanding to a wider range of graph tasks, adding new datasets, etc. Ultimately, we aim to develop ProG into a robust graph prompt toolbox with advanced features like automated prompt selection.
Appendix A Detailed Description of Methods
End-to-end Grpah Models
  • GCN (Graph Convolutional Network [20]): A method that utilizes convolution operation on the graph to propagate information from a node to its neighboring nodes, enabling the network to learn a representation for each node based on its local neighborhood.
  • GraphSAGE (Graph Sample and AggregatE [12]:) A general inductive learning framework that generates node embeddings by sampling and aggregating features from a node’s local neighborhood.
  • GAT (Graph Attention Networks [48]): A graph neural network (GNN) framework that incorporates the attention mechanism. It assigns varying levels of importance to different nodes during the neighborhood information aggregation process, allowing the model to focus on the most informative parts.
  • GT (Graph Transformer [38]): An adaptation of the neural network architecture that applies the principles of the Transformer model to graph-structured data. It uses masks in the self-attention process to leverage the graph structure and enhance model efficiency.
Pre-training
  • DGI[49]: capitalizes on a self-supervised method for pretraining, which is based on the concept of mutual information (MI). It maximizes the MI between the local augmented instances and the global representation.
  • GraphMAE[13]: employs the masked autoencoder framework to minimize the difference between the original features of the masked nodes and their reconstructed features. This process enhances the model’s ability to understand and embed graph structures effectively.
  • EdgePreGPPT[41]: is proposed in GPPT to calculate the dot product as the link probability of the node pair.
  • EdgePreGprompt[28]: is proposed in Gprompt n, for a set of label-free graphs G, we sample several triplets from each graph to increase the similarity between the contextual subgraphs of linked pair while decreasing unlinked pair.
  • GraphCL[59]: applies different graph augmentations to exploit the structural information on the graphs and aims to maximize the agreement between different augmentations for graph pre-training.
  • SimGRACE[54]: tries to perturb the graph model parameter spaces and narrow down the gap between different perturbations for the same graph.
Graph Prompt Methods
  • GPPT (‘Prompt as token’) [41]: introduces a graph prompting function that transforms a standalone node into a token pair composed of a task token, representing a node label, and a structure token, describing the given node. This reformulates the downstream node classification task to resemble edge prediction. As a result, the pre-trained GNNs can be directly applied without extensive fine-tuning to evaluate the linking probability of the token pair and produce the node classification decision. Since GPPT originally is designed for node classification, in this work, to adapt to graph classification, we follow the core thoughts of GPPT and make the following modifications: (1) Introduce $m$ graph task tokens, one for each graph label. (2) Extract $N$ structure tokens from a graph with $N$ nodes, the same as the GPPT for node tasks. (3) Combine each structure token with all graph tokens to form $m$ token pairs. (4) Calculate an $m$-dimension link prediction probability vector for the $m$ structure tokens and choose the class with maximum probability as the prediction result. (6) Aggregate the link prediction results using a voting strategy to determine the graph class. (7) Update the learnable tokens based on the prediction loss.
  • Gprompt (‘Prompt as token’) [28]: employs subgraph similarity as a mechanism to unify different pretext and downstream tasks, including link prediction, node classification, and graph classification. A learnable prompt is subsequently tuned for each downstream task.
  • All-in-one (‘Prompt as graph’) [42]: involves freezing the parameters of pre-trained graph models and incorporating learnable graph prompts as insert patterns to original induced graphs on the downstream node classification task.
  • GPF (‘Prompt as token’) [7]: applies a consistent prompt feature to all nodes, regardless of their contextual differences.
  • GPF-plus (‘Prompt as token’) [7]: enhances the method by assigning prompt features based on node type, ensuring uniformity within each category but without considering the context of individual graphs. Note that GPF and GPF-plus inherently are primarily evaluated in graph classification tasks and are not readily adaptable to node classification. In this work, we address this limitation by inducing a subgraph for each node within a node-level graph. This approach effectively reformulates the node classification task into a graph classification task, allowing for the integration of inserted prompt features.
Appendix B Related Work
B.1 Graph pre-training
Pre-training of graphs is a crucial step in graph representation learning for pre-training, prompting, and predicting paradigms, which provides a solid foundation for downstream tasks by utilizing existing information to encode graph structures. it aims to learn some general knowledge for the graph model with easily accessible information to reduce the annotation costs of new tasks. Node-level pre-training obtains local structural representations, which can be used for downstream tasks. Comparative learning learns by maximizing the mutual information(MI) between the original view and the self-supervised view [5, 62, 18] The prediction model reconstructs perturbed data [12, 13]. However, predictive models are difficult to capture higher-order information. Edge-level pre-training includes distinguishing whether there are edges between node pairs [19, 29], or reconstructing masked edges by restoring including graph reconstruction [34, 55] and contrastive learning targets either local patches of node features and global graph features [49, 40]. Multi-task pre-training addresses multiple optimization objectives, covering various graph-related aspects. This approach enhances generalization and mitigates negative transfer issues. For instance, Hu et al. [15] pre-trained a GNN at both the node and graph levels, allowing the GNN to learn valuable local and global representations simultaneously. Additionally, some studies have utilized contrastive learning at different levels [17, 56], or jointly optimized both contrastive loss and graph reconstruction error [25].
B.2 Graph prompt learning
While Pre-training effectively encodes global information and generates valuable graph-level representations, a significant challenge lies in transferring knowledge from a specific pretext task to downstream tasks with substantial gaps, potentially resulting in negative transfer [35]. Instead of fine-tuning a pre-trained model with an adaptive task head, prompt learning aims to reformulate input data to fit the pretext. Many effective prompt methods were initially developed in the field of NLP, encompassing various approaches such as hand-crafted prompts exemplified by GPT-4 [1], discrete prompts like AutoPrompt [26], and trainable prompts in continuous spaces such as Prefix-Tuning [39]. Due to their remarkable success, prompt-based methods have increasingly been applied to the graph domain.
Graph prompt learning usually has two formalities: on the one hand, it can be regarded as tokens [3, 46, 11]. For example, GPF takes a prompt token as some additional features are added to the original graph. Based on that, GPF-plus [7] trains several independent basis vectors and utilizes attentive aggregation of these basis vectors with the assistance of several learnable linear projections. It also solved problems when training graphs have different scales and large-scale input graphs. GPPT [41] defines graph prompts as additional tokens that contain task tokens and structure tokens that align downstream node task and link prediction pretext. Furthermore, some works [63] use graph prompt to assist graph pooling operation (a.k.a Readout), like Gprompt [28] inserts the prompt vector to the graph pooling by element-wise multiplication. On the other hand, graph prompt is regarded as graphs [10, 16], Such as All-in-one [42] inserts the token graphs as prompt to the original graph and all links between tokens and original graphs, these links can be pre-calculated by the dot product between one token to another token (inner links) or one token to another original node (cross links).
Appendix C Detailed Description of Datasets
Homophilic: Cora and Citeseer datasets [37] consist of a diverse collection of computer science publications, where each node is characterized by a bag-of-words representation of papers and a categorical label indicating the paper topic. Pubmed dataset [30] comprises articles related to diabetes from the PubMed database. Each node in this dataset is represented by an attribute vector containing TF/IDF-weighted word frequencies, accompanied by a label specifying the particular type of diabetes discussed in the publication. ogbn-arxiv [14] is a large-scale directed graph representing the citation network between all Computer Science (CS) arXiv papers indexed by MAG [50]. Each node is an arXiv paper and each directed edge indicates that one paper cites another one. Each paper comes with a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract.
Heterophilic: Cornell, Texas, and Wisconsin are three subdatasets derived from the WebKB dataset [32], compiled from multiple universities’ web pages. Each node within these datasets represents a web page, with edges denoting hyperlinks between pages. The node features are represented as bag-of-words representations of the web pages. Additionally, the web pages are manually categorized into five distinct labels: student, project, course, staff, and faculty. Actor is a co-occurrence network in which each node corresponds to an actor and edges indicate their co-occurrence on the same Wikipedia page [47, 27].
Social networks: IMDB-BINARY [57] is a movie collaboration dataset. The nodes represent actors/actresses, the edges represent collaborations between actors/actresses appearing in the same movie, and the graph labels indicate the genre (Action or Romance). COLLAB [57] is a scientific collaboration dataset. The nodes represent researchers, the edges represent the collaboration relationships between researchers, and the graph labels indicate the field of research (High Energy Physics, Condensed Matter Physics, or Astro Physics).
Biological: PROTEINS [52] is a collection of protein graphs that include the amino acid sequence, conformation, structure, and features such as active sites of the proteins. The nodes represent the secondary structures, and each edge depicts the neighboring relation in the amino-acid sequence or in 3D space. The nodes belong to three categories, and the graphs belong to two classes. ENZYMES [2] is a dataset of 600 enzymes collected from the BRENDA enzyme database. These enzymes are labeled into 6 categories according to their top-level EC enzyme. D&D [6] consists of graph representations of 1,178 proteins. In each graph, nodes represent amino acids, and there is an edge if they are less than six Angstroms apart. Graphs are labeled according to whether they are enzymes or not.
Small Molecule: COX2 [36] is a dataset of molecular structures including 467 cyclooxygenase-2 inhibitors, in which each node is an atom, and each edge represents the chemical bond between atoms, such as single, double, triple or aromatic. All the molecules belong to two categories. BZR [36] is a collection of 405 ligands for benzodiazepine receptors, in which each ligand is represented by a graph. All these ligands belong to 2 categories. MUTAG [21] is a dataset of 188 mutagenic aromatic and heteroaromatic nitro compounds with 7 discrete labels.
Appendix D Additional Details
Metrics. In node and graph classification, AUROC and F1-score are another two important evaluation metrics. AUROC stands for Area Under the Receiver Operating Characteristic Curve. It shows how well the model can distinguish between different classes. A score of 1 means perfect classification, while 0.5 indicates the model is guessing randomly. F1-score combines precision and recall into a single number. Precision measures how many of the predicted positives are actually correct, and recall measures how many of the actual positives the model correctly identified. The F1-score ranges from 0 to 1, with 1 being the best. It is particularly useful in cases where class distributions are imbalanced. In this work, we compute the macro average values for each class. For both AUROC and F1-score, in multi-class classification tasks, each class is treated as the positive class once, while the other classes are treated as negative, and the metrics are averaged.
Implementation Details. To ensure a comprehensive evaluation and maintain fairness across a broad spectrum of models, we develop an open-source package named ProG [^2] https://github.com/sheldonresearch/ProG.
Software and Hardware. For a fair comparison, all methods are implemented by our developed library ProG, which is built upon PyTorch [31], and unless specifically indicated, the encoders for all methods are Graph Convolutional Networks. All experiments are conducted on a Linux server with GPU (NVIDIA GeForce 3090, NVIDIA GeForce 4090, and NVIDIA A800) and CPU (AMD EPYC 7763), using PyTorch 2.0.1 [31], PyG 2.3.1 [8] and Python 3.9.17.
Hyperparameter Settings. To ensure a comprehensive evaluation process and maintain fairness, we conduct an in-depth analysis of various graph prompt methods. Then, we carefully design a unified random search strategy utilized for selecting a set of optimized hyperparameters for each prompt method on each dataset, going beyond the default settings. Given the significant variations in hyperparameters across different prompt methods, we focus on optimizing three types of general hyperparameters via random search. Concretely, we search for the optimal learning rate within the range $10^{uniform(-3,-1)}$, and optimal weight decay within $10^{uniform(-5,-6)}$. Additionally, 32, 64, or 128 batch sizes are selected with equal probability in each trial. Moreover, supervised methods and ’Pre-train & Fine-tune’ methods in this work also undergo the same random search process for hyperparameter optimization.
Appendix E Additional Experimental Analysis
E.1 Adaptivity of Graph Prompts on Different Graphs.
Various Domains. GPL methods, particularly those of the Prompt as Token type, exhibit performance inconsistencies across graph-level datasets from various domains, w.r.t Table 3. Sometimes, these methods may encounter negative transfer issues to varying degrees, except for the BZR and DD datasets. GPPT, in particular, faces significant challenges, likely because first-order subgraphs fail to comprehensively capture graph patterns, introducing noise and affecting transferability.
Large-scale Datasets. Referring to Table 2, on the large-scale node-level dataset (i.e., ogbn-arxiv), Gprompt, GPF, and GPF-plus significantly surpass the Supervised and classical ‘Pre-train & Fine-tune’ methods, demonstrating good scalability. However, the scalability of GPL methods involving subgraph operations, such as GPPT and All-in-one, needs improvement as they fail to fully capture large-scale graph patterns with their subgraph information. Furthermore, regarding the results of the relatively large graph-level dataset (i.e., DD) in Table 3, all GPL methods achieve positive transfer and outperform classical ‘Pre-train & Fine-tune’ methods. Despite this success, the magnitude of improvement indicates that the transferability of GPL methods on large-scale datasets requires further enhancement.
Homophilic & Heterophilic Datasets. Drawing from Table 2, most GPL methods can handle both homophilic and heterophilic graph datasets effectively, with the exception of GPPT. GPPT outperforms the Supervised method on homophilic graph datasets but falls short compared to classical ‘Pre-train & Fine-tune’ methods. On heterophilic datasets, where node attributes differ significantly, GPPT experiences negative transfer issues, possibly due to instability caused by its clustering algorithms.
Pure Structural Dataset. Observing Table 3, the prompting capability for purely structural graph datasets, such as IMDB-B and COLLAB, still requires improvement. While classical ‘Pre-train & Fine-tune’ methods achieve positive transfer on these datasets, some variants of the All-in-one method exhibit negative transfer, highlighting the need for further refinement in handling purely structural graphs.
E.2 Additional Experimental Results with Different Shots and Metrics
The optimal performance of the graph prompt method, along with other methods, under 3-shot and 5-shot learning settings, is displayed in Tables 7- 10. Building on previous experimental outcomes, it is evident that the graph prompt method significantly boosts the effectiveness of various pre-trained graph models in 1/3/5-shot learning settings, thereby enhancing the transferability of pre-trained knowledge in the original ’Pre-train & Fine-tune’ method. Specifically, GPF-plus and All-in-one approaches consistently prove to be the most compatible with different pre-trained methods for node and graph tasks, respectively.
Moreover, the experimental results of various methods under 3/5-shot learning settings for three key metrics—Accuracy, F1-score, and AUROC—are comprehensively presented in Tables 11- 28.
E.3 Curve of The Shot Number and Accuracy
Figure 5 shows the performance changes of different graph prompt methods with increasing shot numbers on node-level dataset Cora and graph-level dataset ENZYMES. The rest baselines are omitted due to their unsatisfactory performance and unstable training process. Specifically, we observe that the performance of various graph prompt methods generally decreases with the increase in shot numbers and may even be surpassed by Supervised methods. This experimental result aligns with the expectation that graph prompt methods are suited for few-shot learning settings because as the number of samples increases and the sample distribution becomes more comprehensive, the effectiveness of graph prompt methods gradually diminishes.
E.4 Performance with More Kinds of Graph Models.
Table 29 and Table 30 show the performance of various graph prompt methods based on different graph models on the Wisconsin and PROTEINS datasets. Overall, compared to the ‘Pre-train & Fine-tune’ method, graph prompt methods are still generally effective across various pre-trained graph models. Among these, the improvement is relatively notable for GAT. Introducing graph prompt techniques to the pre-trained GAT significantly enhances the transferability of the pre-trained model, especially on node-level datasets.
E.5 Other Heat Maps of Other Graph Prompt Methods
Figure 6 displays the heat maps for various graph prompt methods applied to both node-level and graph-level tasks, offering a more intuitive visualization of the enhancements provided by graph prompt methods over the ’Pre-train & Fine-tune’ approach. Through an analysis of these heat maps, it becomes evident that, irrespective of the specific graph prompt method or the type of task, node-level pre-training methods consistently outperform in node-level tasks, whereas graph-level pre-training methods exhibit superior performance in graph-level tasks. Moreover, the experimental results provide further evidence of the robust capability of graph prompt methods in effectively addressing the Negative transfer problem.
Table 7: Comparison of different methods across various datasets on 3-shot node classification tasks. The best results for each dataset are highlighted in bold with a dark red background, and the second-best results are underlined with a light red background.
[…]
Table 8: Comparison of different methods across various datasets on 3-shot graph classification tasks.
[…]
Table 9: Comparison of different methods across various datasets on 5-shot node classification tasks.
[…]
Table 10: Comparison of different methods across various datasets on 5-shot graph classification tasks.
[…]
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 26.56 ±5.55 21.78 ±7.32 39.37 ±16.34 41.60 ±3.10 37.97 ±5.80 20.57 ±4.47 10.99 ±3.19
Pretrain + Fine-tuning DGI 33.15 ±7.84 21.64 ±3.92 42.01 ±12.54 37.49 ±5.13 45.31 ±5.01 19.76 ±3.53 7.21 ±2.91
GraphMAE 32.93 ±3.17 21.26 ±3.57 42.99 ±14.25 36.80 ±7.17 37.81 ±8.62 19.86 ±2.70 12.35 ±3.60
EdgePreGPPT 38.12 ±5.29 18.09 ±5.39 46.74 ±14.09 35.31 ±9.31 47.66 ±2.37 19.17 ±2.53 16.21 ±3.82
EdgePreGprompt 35.57 ±5.83 22.28 ±3.80 41.50 ±7.54 40.69 ±4.13 40.62 ±7.95 20.74 ±4.12 14.83 ±2.38
GCL 52.61 ±1.73 27.02 ±4.31 42.49 ±11.29 33.94 ±7.74 40.31 ±13.68 20.19 ±1.98 4.65 ±1.19
SimGRACE 40.40 ±4.66 35.05 ±4.37 37.59 ±8.17 37.37 ±3.68 46.88 ±4.64 19.78 ±1.89 8.13 ±3.26
GPPTPrompt DGI 30.47 ±3.53 37.26 ±6.17 35.62 ±7.74 29.94 ±10.40 29.29 ±14.57 21.76 ±2.00 3.80 ±6.19
GraphMAE 27.39 ±10.26 21.54 ±4.00 48.31 ±17.72 29.83 ±9.34 25.04 ±10.38 22.58 ±1.97 4.35 ±6.01
EdgePreGPPT 30.37 ±4.30 21.06 ±4.37 39.64 ±7.64 23.89 ±5.40 30.39 ±8.96 19.85 ±0.76 14.65 ±3.07
EdgePreGprompt 25.52 ±4.42 21.85 ±4.30 46.20 ±10.76 30.40 ±6.81 22.68 ±12.82 21.52 ±1.13 2.05 ±1.43
GCL 43.15 ±9.44 26.73 ±4.12 38.34 ±11.59 25.03 ±5.37 31.81 ±15.33 22.51 ±1.73 7.15 ±4.12
SimGRACE 27.86 ±2.79 25.06 ±4.90 36.70 ±9.26 29.83 ±6.44 25.67 ±8.01 20.97 ±2.30 5.50 ±5.10
ALL-in-one DGI 47.52 ±2.50 39.37 ±3.12 38.74 ±2.15 69.54 ±8.98 57.38 ±12.82 21.03 ±1.96 1.93 ±1.48
GraphMAE 23.09 ±4.92 18.08 ±5.23 33.19 ±11.98 67.44 ±10.22 52.82 ±11.47 23.31 ±2.01 4.19 ±6.04
EdgePreGPPT 49.63 ±5.26 35.06 ±2.37 40.73 ±11.32 78.24 ±16.68 58.62 ±5.54 21.49 ±1.27 5.62 ±3.95
EdgePreGprompt 37.39 ±3.31 28.85 ±4.32 35.53 ±9.07 71.40 ±12.32 39.71 ±25.31 20.49 ±3.90 13.01 ±6.29
GCL 52.39 ±10.17 37.37 ±4.15 45.17 ±6.93 42.19 ±1.05 65.49 ±7.06 24.61 ±2.80 6.70 ±6.01
SimGRACE 35.99 ±2.76 40.41 ±2.80 30.23 ±7.03 67.44 ±8.42 59.22 ±20.17 21.03 ±2.23 5.72 ±1.61
Gprompt DGI 36.46 ±5.39 36.25 ±10.26 33.65 ±5.29 74.00 ±4.40 31.00 ±37.32 23.85 ±3.52 48.47 ±5.87
GraphMAE 50.58 ±7.34 42.84 ±10.59 39.74 ±15.35 78.51 ±9.67 23.31 ±29.49 22.34 ±3.57 $6$5.04
EdgePreGPPT 46.96 ±6.22 40.15 ±7.04 35.46 ±14.12 83.80 ±2.44 30.52 ±36.73 23.50 ±4.16 75.72 ±4.95
EdgePreGprompt 48.11 ±9.89 48.07 ±5.62 33.54 ±16.66 81.93 ±2.61 33.25 ±40.11 19.89 ±1.38 70.55 ±7.66
GCL 56.66 ±11.22 45.81 ±7.04 39.37 ±14.95 82.16 ±6.25 29.55 ±35.56 25.26 ±1.10 51.20 ±6.40
SimGRACE 46.34 ±6.75 53.21 ±10.94 35.58 ±9.03 68.95 ±10.71 30.20 ±36.49 24.49 ±4.38 52.76 ±5.30
GPF DGI 27.83 ±18.89 16.50 ±4.57 38.33 ±8.13 74.13 ±1.88 78.45 ±2.57 28.17 ±4.81 17.85 ±8.23
GraphMAE 38.57 ±5.41 25.61 ±3.27 48.52 ±13.23 88.33 ±5.48 78.91 ±18.31 38.37 ±5.82 48.00 ±3.95
EdgePreGPPT 15.29 ±8.41 12.33 ±5.33 43.78 ±6.02 88.67 ±5.78 86.15 ±5.31 25.66 ±3.33 78.37 ±5.47
EdgePreGprompt 26.60 ±13.92 31.16 ±8.05 48.98 ±11.57 82.04 ±4.20 86.26 ±2.52 25.27 ±5.65 54.27 ±11.49
GCL 23.16 ±5.11 16.77 ±1.39 49.99 ±8.86 59.50 ±13.30 87.40 ±3.40 20.68 ±6.70 23.55 ±0.45
SimGRACE 32.01 ±11.21 19.43 ±2.10 37.27 ±6.09 79.48 ±2.62 87.39 ±1.64 28.70 ±3.35 37.56 ±6.10
GPF-plus DGI 17.29 ±6.18 26.60 ±13.24 34.02 ±11.94 76.34 ±6.66 79.54 ±1.89 22.42 ±9.66 16.83 ±10.02
GraphMAE 54.26 ±7.48 59.67 ±11.87 46.64 ±18.57 89.17 ±9.79 94.09 ±4.40 26.58 ±7.84 49.81 ±2.62
EdgePreGPPT 28.49 ±18.73 28.04 ±14.31 46.51 ±15.84 91.03 ±4.11 92.14 ±5.46 29.32 ±8.56 71.98 ±12.23
EdgePreGprompt 55.77 ±10.30 49.43 ±8.21 42.79 ±18.18 85.44 ±2.23 87.70 ±3.77 22.68 ±3.64 57.44 ±6.95
GCL 34.18 ±17.71 28.86 ±22.88 37.02 ±11.29 68.57 ±8.77 95.83 ±4.19 22.82 ±4.99 32.11 ±4.86
SimGRACE 21.33 ±14.86 24.61 ±21.21 35.90 ±9.06 73.64 ±5.25 90.74 ±3.12 20.51 ±4.24 46.71 ±3.17
Table 12: Node classification performance F1-score (1-shot)
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 16.60±2.54 10.81±4.90 37.23±15.48 26.34±4.01 24.05±5.12 11.56±3.08 7.99±0.99
Pretrain + Fine-tuning DGI 24.96±6.01 11.01±5.94 34.75±13.75 26.69±3.39 28.90±6.81 12.01±1.43 3.93±1.32
GraphMAE 23.18±2.85 10.82±3.83 41.03±13.36 27.43±4.47 23.08±6.07 12.71±1.24 8.07±1.08
EdgePreGPPT 35.92±4.06 10.86±2.29 40.62±18.09 23.56±3.43 29.03±5.16 14.59±3.09 12.13±2.04
EdgePreGprompt 28.99±6.35 12.39±3.56 28.89±6.74 26.74±3.28 26.81±6.66 11.63±2.64 10.61±1.45
GCL 47.14±3.15 21.86±3.67 38.30±10.89 14.27±3.90 24.52±7.59 15.91±0.98 2.88±0.51
SimGRACE 33.22±3.29 30.78±3.91 32.79±6.60 23.71±2.97 29.53±6.44 15.73±1.20 2.88±0.50
GPPTPrompt DGI 25.82±4.78 33.00±6.49 31.92±8.94 22.77±6.29 23.80±8.89 17.59±1.13 0.17±0.26
GraphMAE 13.18±6.02 10.87±4.43 46.43±16.73 23.74±5.95 21.36±7.80 13.60±1.69 0.27±0.23
EdgePreGPPT 28.54±3.87 17.62±4.24 34.55±8.09 20.61±5.49 23.84±5.22 18.48±0.57 9.15±1.18
EdgePreGprompt 23.46±5.11 19.00±4.24 45.52±10.54 23.65±3.96 19.82±7.93 19.39±1.08 1.35±1.07
GCL 38.99±8.32 23.76±3.97 36.75±12.92 18.08±6.69 25.64±8.12 19.62±0.56 1.52±0.68
SimGRACE 21.70±2.66 22.13±4.48 31.55±10.44 21.48±3.16 21.75±5.75 17.15±1.75 0.66±0.39
ALL-in-one DGI 34.76±3.89 26.67±4.46 22.68±5.29 52.08±6.70 31.14±20.37 14.37±2.36 0.24±0.21
GraphMAE 10.82±4.45 7.04±2.29 25.46±9.17 53.76±2.54 37.08±9.37 12.68±2.63 0.19±0.26
EdgePreGPPT 44.16±4.02 26.79±3.27 36.27±12.89 67.68±10.36 26.56±10.81 14.72±3.29 0.98±0.47
EdgePreGprompt 24.93±4.99 14.58±5.03 30.99±7.62 62.98±8.70 28.81±17.76 11.85±1.48 0.56±0.27
GCL 46.58±8.42 29.35±3.66 38.05±6.24 34.44±5.73 43.37±16.01 16.05±3.88 0.75±1.07
SimGRACE 27.35±1.31 30.20±4.44 24.61±6.29 52.59±6.43 39.01±18.76 13.86±1.76 1.28±0.35
Gprompt DGI 30.20±4.21 33.58±9.40 31.89±5.43 61.63±9.08 23.68±28.78 20.12±3.96 42.90±1.91
GraphMAE 45.91±6.10 40.94±10.71 39.46±15.97 68.64±8.14 19.27±23.52 20.06±4.69 57.88±3.32
EdgePreGPPT 44.15±7.57 37.31±6.26 29.87±12.41 74.76±4.26 24.14±29.42 20.14±3.34 69.86±1.92
EdgePreGprompt 46.28±8.46 42.82±6.05 33.57±16.07 77.03±6.40 25.86±31.46 18.67±1.74 69.89±5.15
GCL 49.86±10.36 40.41±9.12 38.04±13.45 65.53±6.42 21.15±25.84 22.00±1.74 45.32±3.89
SimGRACE 38.55±5.02 49.65±11.42 31.13±9.15 61.22±8.58 29.20±35.62 21.39±3.95 46.73±4.62
GPF DGI 19.39±17.53 4.68±1.17 28.70±9.13 57.50±2.20 68.55±8.62 19.25±10.07 6.59±3.79
GraphMAE 23.79±5.49 12.89±1.99 45.36±15.88 82.97±6.10 73.83±15.04 31.69±5.47 19.47±3.26
EdgePreGPPT 3.76±1.75 3.59±1.38 31.20±15.26 78.86±7.05 72.46±12.05 18.04±8.30 63.46±4.32
EdgePreGprompt 15.85±13.63 18.63±7.34 39.90±9.00 80.44±7.43 78.43±9.49 19.72±8.08 29.38±6.81
GCL 9.76±4.57 4.78±0.34 38.27±17.17 35.78±8.46 76.01±7.84 11.69±9.39 18.18±1.66
SimGRACE 18.73±10.08 5.90±1.13 22.65±5.69 70.22±3.85 74.43±7.73 17.94±5.36 14.16±1.23
GPF-plus DGI 9.94±8.40 13.74±16.38 24.21±9.44 62.02±12.59 67.91±7.85 14.51±10.72 5.37±4.20
GraphMAE 50.19±10.64 56.22±13.99 42.38±19.01 84.72±15.08 81.36±10.39 19.25±5.53 23.38±1.12
EdgePreGPPT 19.52±17.74 15.32±19.12 40.64±17.93 85.24±5.45 78.97±3.30 24.56±8.79 60.35±7.73
EdgePreGprompt 53.28±11.46 48.37±9.28 40.02±18.18 74.18±6.55 75.46±5.42 15.58±4.14 31.37±2.50
GCL 26.14±21.11 18.90±24.01 33.64±11.41 47.94±17.19 86.22±10.29 16.08±3.05 15.78±1.60
SimGRACE 14.35±19.65 15.23±20.93 23.74±12.56 52.05±5.74 80.81±8.01 11.83±4.68 21.81±1.06
Table 13: Node classification performance AUROC (1-shot)
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 76.71±1.76 63.02±4.09 62.01±18.06 61.53±4.96 62.86±10.22 51.16±1.01 72.48±1.34
Pretrain + Fine-tuning DGI 79.33±2.82 65.16±4.85 63.47±13.36 59.72±2.84 65.16±9.96 50.13±0.96 64.22±2.71
GraphMAE 78.93±0.67 63.92±3.46 66.78±16.67 60.83±5.18 61.79±9.63 51.01±0.55 72.86±1.61
EdgePreGPPT 78.44±1.58 60.74±3.98 63.49±15.75 58.91±3.16 65.05±9.01 50.79±0.65 77.26±2.75
EdgePreGprompt 80.12±2.93 65.05±4.65 60.89±13.56 61.08±5.22 62.77±10.98 51.47±1.28 75.81±2.26
GCL 86.34±1.24 68.32±1.24 57.56±15.85 43.79±4.25 63.09±10.20 51.19±0.63 63.36±1.49
SimGRACE 76.63±1.67 71.91±2.47 58.10±9.13 57.76±2.12 65.06±9.74 50.81±0.66 56.33±0.77
GPPTPrompt DGI 63.97 ±5.40 71.52 ±5.15 51.57 ±5.86 59.89 ±2.95 57.22 ±9.90 50.71 ±0.86 50.07 ±0.08
GraphMAE 78.20 ±3.89 61.64 ±4.31 64.12 ±15.44 55.48 ±6.44 53.95 ±3.70 51.71 ±1.74 59.28 ±2.39
EdgePreGPPT 65.75 ±4.04 54.28 ±5.62 55.12 ±11.88 54.59 ±4.65 50.55 ±3.83 49.94 ±0.53 74.38 ±1.59
EdgePreGprompt 62.60 ±2.50 54.23 ±6.14 61.41 ±11.42 57.58 ±2.23 53.16 ±6.75 50.32 ±0.29 63.71 ±2.13
GCL 72.09 ±7.61 59.06 ±3.51 55.95 ±13.50 54.51 ±4.05 56.59 ±7.31 51.76 ±0.95 54.45 ±1.08
SimGRACE 64.83 ±2.80 59.61 ±4.19 51.84 ±12.98 56.88 ±4.39 53.42 ±6.67 50.48 ±0.77 50.72 ±0.43
ALL-in-one DGI 75.29 ±1.07 71.50 ±1.46 55.69 ±1.38 78.94 ±2.46 64.75 ±6.31 51.05 ±0.96 55.21 ±1.45
GraphMAE 73.99 ±5.63 49.98 ±3.07 53.18 ±13.11 81.15 ±2.57 62.90 ±3.67 50.96 ±1.97 50.21 ±1.01
EdgePreGPPT 83.26 ±1.22 69.82 ±1.32 73.63 ±3.06 87.52 ±6.90 75.38 ±10.13 49.99 ±0.28 54.22 ±5.18
EdgePreGprompt 73.84 ±3.01 62.35 ±6.11 61.26 ±7.66 89.84 ±3.19 67.99 ±6.33 51.48 ±2.87 75.81 ±2.83
GCL 84.34 ±4.32 72.36 ±2.82 68.22 ±6.38 72.40 ±3.63 69.39 ±2.56 47.63 ±2.10 55.77 ±4.78
SimGRACE 72.30±1.28 69.88±0.83 52.94±2.50 81.86±5.69 70.92±7.19 48.18±0.97 65.43±1.59
Gprompt DGI 70.70 ±2.26 70.93 ±4.80 52.96 ±2.53 85.91 ±2.99 57.93 ±10.19 52.82 ±3.88 90.60 ±1.08
GraphMAE 80.67 ±5.24 70.72 ±5.34 60.66 ±19.63 93.63 ±2.61 59.57 ±11.81 52.87 ±4.78 94.04 ±1.32
EdgePreGPPT 84.03 ±2.26 67.95 ±2.69 44.60 ±13.08 95.25 ±2.14 60.65 ±12.95 53.55 ±3.61 96.40 ±0.74
EdgePreGprompt 81.90 ±4.04 74.85 ±2.68 58.34 ±22.51 96.17 ±1.26 62.03 ±14.44 50.16 ±1.75 94.39 ±1.12
GCL 83.03 ±4.16 78.33 ±5.28 58.24 ±13.76 88.03 ±5.49 58.33 ±10.63 54.45 ±2.93 92.72 ±0.72
SimGRACE 76.99 ±3.17 84.09 ±2.88 51.91 ±12.26 86.09 ±6.13 61.60 ±14.00 53.63 ±4.60 91.93 ±1.25
GPF DGI 65.74 ±10.40 50.00 ±0.00 51.90 ±7.57 83.72 ±4.30 68.70 ±1.83 58.84 ±3.16 64.74 ±2.99
GraphMAE 73.69 ±2.75 66.16 ±0.22 73.35 ±8.02 98.84 ±0.77 84.20 ±10.56 65.43 ±3.58 77.64 ±0.97
EdgePreGPPT 62.03 ±6.87 57.99 ±6.03 68.25 ±7.09 98.54 ±0.99 84.45 ±7.94 55.90 ±6.20 93.53 ±1.06
EdgePreGprompt 70.83 ±8.42 71.94 ±3.31 71.40 ±8.27 97.48 ±0.73 81.31 ±7.35 56.13 ±4.33 82.60 ±2.47
GCL 74.17 ±5.66 61.21 ±7.77 70.39 ±4.76 62.05 ±4.35 82.48 ±5.23 54.42 ±3.73 73.36 ±1.69
SimGRACE 69.60 ±10.54 62.47 ±3.70 50.73 ±3.64 93.71 ±1.37 85.70 ±8.79 60.48 ±2.78 61.29 ±1.33
GPF-plus DGI 58.07 ±8.25 55.15 ±11.69 55.70 ±11.04 91.39 ±4.68 78.71 ±9.15 57.02 ±8.99 60.73 ±3.11
GraphMAE 86.94 ±4.26 87.43 ±3.46 66.74 ±18.62 99.25 ±1.27 86.83 ±8.56 63.32 ±6.43 80.32 ±3.00
EdgePreGPPT 68.27 ±10.69 59.85 ±17.18 71.27 ±10.40 99.12 ±1.13 83.10 ±7.26 61.72 ±5.16 93.52 ±1.09
EdgePreGprompt 86.33 ±4.94 88.60 ±5.18 63.12 ±18.73 96.59 ±3.43 82.30 ±6.38 61.97 ±5.10 84.36 ±0.87
GCL 80.06 ±6.24 61.58 ±14.77 55.63 ±15.68 77.34 ±13.37 83.65 ±8.19 58.86 ±3.80 70.39 ±3.22
SimGRACE 64.39 ±11.25 64.11 ±10.53 54.79 ±10.68 88.74 ±2.84 86.70 ±8.72 55.75 ±4.51 69.09 ±2.34
[…]
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 37.79 ±9.16 35.18 ±6.86 57.33 ±4.64 41.03 ±6.40 40.78 ±12.55 18.62 ±3.46 19.03 ±5.08
Pretrain + Fine-tuning DGI 45.84 ±3.29 34.91 ±10.07 63.00 ±6.83 39.43 ±6.93 43.13 ±13.79 20.27 ±1.70 21.85 ±5.37
GraphMAE 45.15 ±4.77 28.59 ±6.41 65.40 ±3.00 41.49 ±5.19 40.94 ±13.91 19.03 ±3.17 19.63 ±7.64
EdgePreGPPT 51.97 ±2.84 33.19 ±6.79 64.38 ±3.59 42.40 ±7.77 34.69 ±12.07 20.83 ±2.24 20.51 ±4.02
EdgePredGprompt 40.33 ±6.62 33.39 ±5.51 63.49 ±3.17 41.37 ±7.28 34.53 ±10.09 19.29 ±1.95 27.34 ±6.61
GCL 49.39 ±9.15 38.40 ±3.06 62.79 ±3.21 41.26 ±6.26 40.31 ±13.36 21.06 ±1.27 17.37 ±7.34
SimGRACE 43.61 ±5.41 45.08 ±2.09 62.66 ±3.21 40.11 ±8.04 40.94 ±13.73 22.11 ±1.97 16.06 ±2.05
GPPTPrompt DGI 37.46 ±6.27 42.34 ±8.31 44.36 ±3.67 34.29 ±4.71 37.80 ±8.98 21.42 ±3.13 14.34 ±14.10
GraphMAE 30.93 ±3.64 20.76 ±2.64 67.43 ±2.96 33.60 ±2.30 37.64 ±5.88 21.65 ±3.39 17.24 ±13.25
EdgePreGPPT 35.05 ±2.95 24.26 ±2.55 58.66 ±5.93 29.37 ±6.22 38.90 ±7.56 20.35 ±0.43 22.46 ±4.05
EdgePreGprompt 27.94 ±5.07 23.21 ±2.95 64.98 ±5.35 33.49 ±5.49 36.06 ±12.01 19.85 ±0.31 18.73 ±4.77
GCL 43.84 ±6.11 27.09 ±4.57 51.30 ±6.35 23.89 ±5.35 38.90 ±8.86 20.60 ±1.10 13.04 ±6.44
SimGRACE 29.66 ±5.49 29.63 ±1.87 43.95 ±3.55 31.66 ±6.21 34.65 ±6.97 21.08 ±1.16 14.13 ±13.89
ALL-in-one DGI 33.80 ±6.36 31.08 ±2.95 48.09 ±2.81 74.71 ±2.35 31.94 ±7.00 24.23 ±1.39 29.92 ±1.52
GraphMAE 19.25 ±3.11 19.02 ±5.19 64.03 ±4.70 76.90 ±2.98 65.93 ±17.42 19.56 ±2.48 13.15 ±4.63
EdgePreGPPT 48.09 ±4.83 34.29 ±1.85 58.36 ±4.63 89.62 ±4.38 88.69 ±1.08 20.91 ±3.37 14.41 ±5.57
EdgePreGprompt 29.54 ±6.25 24.63 ±3.57 65.79 ±5.79 84.29 ±2.32 73.89 ±16.92 20.70 ±2.58 12.91 ±2.08
GCL 27.89 ±18.34 43.96 ±6.69 51.44 ±4.46 67.06 ±9.70 78.86 ±4.06 22.14 ±2.57 31.15 ±2.25
SimGRACE 25.16 ±11.78 48.09 ±8.18 47.87 ±3.53 76.78 ±1.28 82.93 ±5.04 21.99 ±1.17 15.47 ±4.05
Gprompt DGI 37.42 ±7.07 41.19 ±5.30 38.69 ±1.54 71.54 ±10.15 0.50 ±0.41 26.64 ±4.45 19.40 ±5.30
GraphMAE 63.78 ±5.77 57.15 ±3.90 62.47 ±4.29 86.48 ±7.48 0.50 ±0.41 29.21 ±3.27 18.90 ±3.10
EdgePreGPPT 51.29 ±7.07 34.79 ±4.07 53.63 ±7.62 89.10 ±4.09 0.50 ±0.41 26.44 ±3.36 18.14 ±3.54
EdgePreGprompt 53.37 ±6.30 56.19 ±2.95 66.68 ±3.53 92.52 ±5.38 0.50 ±0.41 21.78 ±1.85 19.59 ±3.20
GCL 56.61 ±8.02 56.28 ±5.89 50.33 ±5.18 85.83 ±4.53 0.50 ±0.41 29.67 ±2.53 18.94 ±2.93
SimGRACE 45.22 ±5.69 60.00 ±6.18 43.58 ±5.78 82.84 ±1.52 0.50 ±0.41 25.07 ±1.10 15.74 ±4.18
GPF DGI 25.14 ±19.07 25.92 ±12.30 38.49 ±14.50 76.22 ±5.46 44.94 ±4.84 25.03 ±1.66 29.90 ±3.01
GraphMAE 33.93 ±8.01 23.91 ±5.11 71.20 ±2.82 98.92 ±0.79 94.67 ±3.28 37.44 ±3.43 12.22 ±5.44
EdgePreGPPT 23.05 ±9.03 18.81 ±4.35 65.06 ±2.22 96.63 ±2.53 94.85 ±1.67 30.47 ±3.57 19.76 ±4.75
EdgePreGprompt 32.57 ±9.57 21.75 ±2.88 67.57 ±14.07 94.58 ±5.67 98.49 ±1.45 25.64 ±8.58 15.72 ±5.16
GCL 34.84 ±19.83 23.02 ±5.56 49.15 ±6.46 56.93 ±12.66 94.66 ±3.87 21.95 ±6.54 8.90 ±2.80
SimGRACE 27.80 ±18.09 21.20 ±11.25 31.75 ±13.86 86.10 ±0.76 97.83 ±1.13 22.82 ±3.16 10.71 ±1.11
GPF-plus DGI 20.80 ±13.69 24.46 ±17.23 51.70 ±11.29 85.61 ±8.19 72.61 ±22.35 23.99 ±6.98 16.16 ±6.23
GraphMAE 56.38 ±5.37 72.48 ±5.63 70.85 ±4.03 97.44 ±4.52 97.00 ±0.68 43.59 ±4.52 22.50 ±14.10
EdgePreGPPT 32.62 ±23.21 22.06 ±21.01 67.20 ±1.67 97.34 ±2.33 98.83 ±0.85 34.38 ±5.23 23.56 ±5.67
EdgePreGprompt 52.74 ±7.08 71.56 ±8.03 68.84 ±4.33 96.96 ±4.60 97.01 ±1.62 30.87 ±6.61 24.71 ±6.77
GCL 30.34 ±21.41 27.17 ±24.65 61.23 ±11.39 62.76 ±14.68 97.18 ±1.93 36.87 ±2.45 5.42 ±0.82
SimGRACE 30.95 ±19.32 28.97 ±20.66 41.96 ±15.64 80.64 ±3.99 97.34 ±1.84 34.89 ±6.03 13.16 ±5.89
Table 15: Node classification performance F1-score (3-shot)
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 37.64 ±6.32 28.68 ±8.09 53.04 ±6.01 34.90 ±5.93 29.16 ±7.77 13.69 ±2.87 15.21 ±2.66
Pretrain + Fine-tuning DGI 41.76 ±1.48 27.23 ±9.82 61.18 ±10.18 35.01 ±5.55 31.12 ±8.67 16.05 ±0.75 16.45 ±1.97
GraphMAE 45.70 ±6.35 20.60 ±5.72 64.33 ±3.42 34.08 ±4.01 28.29 ±8.05 16.30 ±2.37 14.26 ±4.46
EdgePreGPPT 49.76 ±2.52 28.14 ±6.04 63.62 ±4.81 35.03 ±5.65 25.55 ±6.58 17.79 ±1.62 16.68 ±2.65
EdgePredGprompt 41.69 ±5.58 23.81 ±4.86 62.48 ±4.02 35.42 ±6.02 26.05 ±6.95 12.98 ±2.02 19.33 ±2.46
GCL 47.79 ±9.36 34.47 ±3.44 63.10 ±2.93 22.42 ±6.61 29.45 ±8.86 19.91 ±1.08 7.46 ±3.10
SimGRACE 41.71 ±4.09 42.01 ±1.56 63.14 ±3.43 34.28 ±5.18 29.55 ±7.96 17.92 ±1.61 11.68 ±1.37
GPPTPrompt DGI 33.94 ±7.56 36.13 ±8.65 44.13 ±3.80 30.91 ±2.41 34.81 ±7.09 19.30 ±2.98 3.50 ±2.51
GraphMAE 25.31 ±6.18 6.91 ±2.93 67.61 ±2.62 29.51 ±3.03 33.87 ±3.38 18.03 ±1.97 6.06 ±5.21
EdgePreGPPT 32.77 ±2.69 19.54 ±1.23 58.07 ±6.41 27.20 ±3.72 33.65 ±5.16 19.32 ±0.53 17.24 ±1.62
EdgePreGprompt 20.10 ±4.36 20.60 ±1.98 62.73 ±4.96 29.19 ±4.29 30.98 ±7.36 18.99 ±0.70 15.47 ±2.95
GCL 41.25 ±4.50 25.71 ±4.25 50.27 ±6.65 15.65 ±5.40 34.61 ±6.65 19.81 ±1.02 5.96 ±3.82
SimGRACE 27.19 ±4.01 27.09 ±2.43 42.79 ±3.17 26.79 ±4.55 32.38 ±5.73 19.02 ±0.50 4.01 ±2.59
ALL-in-one DGI 22.05 ±4.15 25.31 ±3.49 42.82 ±4.66 63.22 ±7.20 37.71 ±11.00 11.04 ±2.36 15.79 ±0.87
GraphMAE 8.10 ±2.43 7.36 ±2.81 62.24 ±5.80 51.62 ±2.63 58.91 ±14.27 11.52 ±0.97 9.72 ±3.20
EdgePreGPPT 41.54 ±1.81 27.24 ±3.58 56.20 ±5.02 80.48 ±6.96 82.89 ±2.60 9.51 ±3.89 9.61 ±4.38
EdgePreGprompt 20.33 ±9.09 12.09 ±3.36 63.18 ±7.52 79.59 ±6.58 68.71 ±14.50 7.89 ±2.38 9.96 ±1.24
GCL 14.85 ±18.09 33.38 ±6.65 50.52 ±3.48 61.79 ±11.38 69.09 ±6.96 16.40 ±1.47 14.33 ±0.41
SimGRACE 12.71 ±11.42 40.22 ±8.61 44.79 ±2.93 64.88 ±4.35 79.46 ±5.10 13.70 ±2.29 10.07 ±1.80
Gprompt DGI 32.92 ±6.82 37.71 ±5.52 37.90 ±1.53 60.12 ±10.43 0.20 ±0.16 26.39 ±4.24 14.59 ±2.78
GraphMAE 59.30 ±7.36 53.60 ±3.06 61.33 ±4.81 79.06 ±11.48 0.20 ±0.16 29.02 ±3.53 14.78 ±1.50
EdgePreGPPT 46.21 ±6.52 29.81 ±3.69 48.94 ±10.06 80.60 ±7.17 0.20 ±0.16 24.73 ±3.32 15.68 ±2.29
EdgePreGprompt 52.43 ±6.41 53.88 ±3.50 65.02 ±2.80 91.03 ±5.30 0.20 ±0.16 20.40 ±1.65 15.56 ±1.69
GCL 51.54 ±8.49 52.46 ±5.08 49.16 ±5.73 65.08 ±11.12 0.20 ±0.16 28.21 ±2.34 14.80 ±1.34
SimGRACE 40.83 ±3.88 56.25 ±7.01 42.59 ±5.51 76.68 ±3.65 0.20 ±0.16 25.41 ±1.35 12.58 ±1.91
GPF DGI 12.93 ±16.42 12.92 ±13.22 32.76 ±17.33 67.09 ±7.72 46.87 ±5.91 15.68 ±9.94 14.93 ±1.60
GraphMAE 19.59 ±5.22 14.48 ±5.04 70.15 ±2.68 94.99 ±4.12 83.16 ±10.27 31.56 ±8.62 5.27 ±5.58
EdgePreGPPT 9.90 ±9.04 7.40 ±3.59 63.04 ±3.09 92.67 ±6.23 80.32 ±6.04 22.38 ±8.76 14.77 ±2.22
EdgePreGprompt 21.59 ±5.95 11.79 ±3.50 62.02 ±21.60 93.30 ±5.43 86.36 ±10.64 22.04 ±11.41 11.51 ±3.42
GCL 18.90 ±20.40 8.10 ±4.95 47.10 ±8.36 30.08 ±6.91 81.66 ±11.83 14.24 ±8.36 4.28 ±1.13
SimGRACE 13.21 ±14.54 10.26 ±8.85 25.30 ±17.26 77.51 ±2.92 85.34 ±10.18 15.37 ±7.26 5.41 ±1.24
GPF-plus DGI 10.87 ±13.23 14.35 ±19.57 43.25 ±20.47 78.25 ±14.06 63.25 ±15.75 14.70 ±3.64 10.41 ±3.06
GraphMAE 55.76 ±3.81 70.45 ±6.54 69.78 ±3.58 97.33 ±2.96 83.31 ±9.41 45.02 ±4.47 12.22 ±4.77
EdgePreGPPT 20.90 ±24.90 15.30 ±23.70 65.43 ±1.28 94.18 ±5.80 86.54 ±10.00 32.74 ±6.14 14.90 ±2.10
EdgePreGprompt 52.13 ±7.14 68.79 ±9.94 67.98 ±3.69 94.78 ±4.60 85.14 ±8.73 30.59 ±7.43 15.43 ±2.65
GCL 21.84 ±23.15 18.50 ±28.28 59.42 ±13.43 47.29 ±15.14 84.24 ±8.20 30.93 ±5.62 4.37 ±1.56
SimGRACE 20.11 ±22.26 18.43 ±24.95 27.58 ±21.15 61.92 ±12.87 85.24 ±9.54 30.11 ±6.73 6.91 ±2.30
Table 16: Node classification performance AUROC (3-shot)
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 86.20 ±2.47 76.31 ±3.45 80.24 ±2.82 66.44 ±6.12 59.85 ±12.05 51.48 ±0.74 59.17 ±1.26
Pretrain + Fine-tuning DGI 84.23 ±2.75 74.32 ±4.21 82.28 ±5.43 67.49 ±3.80 55.28 ±10.97 51.93 ±1.20 61.28 ±1.40
GraphMAE 86.80 ±2.35 75.07 ±2.48 84.70 ±1.28 66.58 ±6.51 58.48 ±11.37 51.78 ±1.01 59.12 ±1.66
EdgePreGPPT 85.62 ±1.41 72.39 ±3.60 83.03 ±2.89 66.70 ±4.04 54.49 ±12.28 52.25 ±0.89 60.55 ±1.73
EdgePredGprompt 85.85 ±0.79 76.73 ±2.26 83.40 ±2.74 67.27 ±4.41 58.43 ±10.76 51.24 ±0.96 61.30 ±1.07
GCL 87.71 ±1.59 75.10 ±2.66 81.83 ±2.15 61.56 ±7.86 57.10 ±12.53 52.03 ±1.09 54.66 ±2.69
SimGRACE 80.39 ±2.84 76.94 ±1.30 79.56 ±3.24 67.34 ±5.82 59.02 ±11.26 51.60 ±0.77 56.28 ±1.14
GPPTPrompt DGI 70.88 ±2.54 73.97 ±5.46 61.07 ±2.93 66.79 ±2.66 63.59 ±7.91 51.34 ±1.54 50.34 ±0.68
GraphMAE 80.91 ±2.38 69.11 ±3.46 83.10 ±1.03 63.77 ±3.61 60.95 ±3.13 51.22 ±1.38 51.84 ±3.86
EdgePreGPPT 68.71 ±2.02 59.01 ±2.73 74.22 ±4.90 57.60 ±1.78 58.20 ±4.72 50.23 ±0.31 58.95 ±1.29
EdgePreGprompt 68.29 ±3.52 58.07 ±2.40 81.15 ±4.81 61.93 ±5.66 61.91 ±5.28 50.27 ±0.63 56.16 ±2.47
GCL 75.26 ±3.86 60.58 ±4.04 71.43 ±2.47 53.79 ±2.21 61.73 ±5.50 50.96 ±0.66 50.69 ±1.68
SimGRACE 66.90 ±4.18 65.72 ±2.09 62.58 ±2.08 66.87 ±2.38 60.35 ±7.62 50.53 ±0.57 50.07 ±0.24
ALL-in-one DGI 71.15 ±3.33 65.99 ±0.90 66.83 ±1.02 81.86 ±0.82 46.05 ±2.66 52.80 ±2.23 55.31 ±0.33
GraphMAE 77.89 ±3.94 63.31 ±2.05 84.13 ±1.99 81.57 ±1.86 66.49 ±5.26 50.76 ±1.17 55.43 ±0.07
EdgePreGPPT 80.90 ±2.72 68.74 ±3.28 76.23 ±0.81 96.17 ±3.82 75.39 ±1.30 52.98 ±1.21 56.37 ±0.72
EdgePreGprompt 77.50 ±2.96 69.84 ±1.80 85.03 ±1.73 91.35 ±2.44 72.04 ±6.05 49.89 ±1.95 54.08 ±0.15
GCL 65.18 ±11.80 77.64 ±2.06 73.05 ±1.97 82.74 ±3.64 72.60 ±1.20 51.67 ±1.31 53.03 ±0.27
SimGRACE 59.44 ±9.08 78.11 ±1.17 64.76 ±2.62 84.99 ±2.40 92.99 ±1.95 51.64 ±2.25 51.35 ±0.09
Gprompt DGI 62.24 ±2.99 75.87 ±3.70 55.46 ±1.89 87.72 ±5.97 49.44 ±8.34 55.78 ±3.19 52.89 ±2.11
GraphMAE 80.27 ±5.61 74.38 ±3.36 84.37 ±1.75 96.34 ±2.81 49.44 ±8.34 58.52 ±2.66 53.56 ±1.76
EdgePreGPPT 81.41 ±5.40 65.99 ±4.06 72.84 ±0.48 98.59 ±1.16 49.44 ±8.34 56.86 ±2.69 56.05 ±3.09
EdgePreGprompt 83.15 ±3.91 82.01 ±1.78 84.45 ±1.80 99.32 ±0.66 49.44 ±8.34 51.33 ±2.32 54.94 ±1.15
GCL 83.25 ±3.22 83.20 ±0.97 71.35 ±3.39 91.20 ±4.59 49.44 ±8.34 59.78 ±3.41 53.58 ±1.71
SimGRACE 79.02 ±2.55 86.57 ±1.86 61.89 ±6.64 94.09 ±3.00 49.44 ±8.34 56.13 ±1.59 52.64 ±1.18
GPF DGI 60.20 ±13.06 59.79 ±13.56 59.52 ±12.99 84.92 ±3.03 71.79 ±11.35 54.39 ±2.39 55.76 ±1.34
GraphMAE 82.70 ±2.62 72.25 ±4.60 87.91 ±1.01 99.59 ±0.81 91.54 ±9.54 67.24 ±4.73 55.48 ±0.92
EdgePreGPPT 65.44 ±9.07 67.13 ±1.95 81.07 ±0.78 99.62 ±0.64 84.29 ±6.85 58.43 ±3.84 58.43 ±1.11
EdgePreGprompt 83.18 ±3.22 69.63 ±7.04 85.38 ±6.75 99.56 ±0.58 87.93 ±6.62 56.19 ±3.18 54.51 ±2.74
GCL 73.76 ±9.73 68.34 ±10.91 75.10 ±2.04 52.13 ±1.68 87.77 ±8.39 56.96 ±4.69 51.48 ±0.80
SimGRACE 63.96 ±7.80 63.26 ±11.65 58.71 ±6.49 95.66 ±0.57 90.93 ±9.06 55.06 ±3.04 49.92 ±0.31
GPF-plus DGI 55.30 ±13.20 55.81 ±15.74 66.54 ±15.31 93.25 ±5.92 77.44 ±7.96 57.09 ±3.17 55.62 ±1.15
GraphMAE 87.26 ±3.14 91.53 ±1.55 87.12 ±1.71 $9$0.02 86.23 ±6.11 72.39 ±1.82 58.01 ±1.44
EdgePreGPPT 73.87 ±10.93 56.16 ±16.39 82.80 ±1.90 99.75 ±0.50 87.25 ±8.99 67.15 ±3.76 57.48 ±1.34
EdgePreGprompt 85.81 ±3.03 91.47 ±1.96 85.13 ±2.57 99.73 ±0.51 88.28 ±9.02 65.17 ±4.06 56.91 ±1.33
GCL 80.93 ±6.55 61.66 ±15.01 80.00 ±7.05 78.70 ±14.44 81.41 ±2.06 69.40 ±1.84 51.15 ±0.58
SimGRACE 66.36 ±16.56 60.91 ±15.05 61.66 ±12.28 90.60 ±4.75 91.16 ±9.14 66.44 ±4.39 49.94 ±1.04
[…]
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 50.25 ±8.37 41.22 ±6.30 67.88 ±2.18 39.43 ±5.86 43.91 ±6.47 21.92 ±1.86 22.38 ±3.05
Pretrain + Fine-tuning DGI 48.79 ±8.51 35.91 ±4.94 61.44 ±7.93 40.00 ±6.31 47.19 ±7.37 21.48 ±1.71 12.48 ±6.15
GraphMAE 54.09 ±4.21 32.75 ±6.35 70.04 ±4.57 40.69 ±6.73 43.44 ±8.48 22.92 ±1.22 24.08 ±1.45
EdgePreGPPT 54.34 ±3.78 38.97 ±7.35 70.91 ±4.87 40.69 ±6.77 44.53 ±11.05 21.46 ±1.00 28.84 ±3.11
EdgePredGprompt 49.04 ±8.99 32.08 ±8.24 70.44 ±5.04 41.03 ±6.58 45.78 ±8.07 22.46 ±1.99 25.94 ±2.63
GCL 62.66 ±3.55 39.54 ±3.54 66.84 ±5.78 42.97 ±8.99 46.41 ±5.47 21.99 ±1.61 13.42 ±3.73
SimGRACE 45.13 ±7.81 38.90 ±6.03 62.65 ±6.29 41.49 ±5.77 46.09 ±8.71 22.77 ±1.56 9.45 ±1.62
GPPTPrompt DGI 43.68 ±7.12 45.77 ±7.41 47.39 ±10.22 36.29 ±3.97 48.82 ±5.15 20.91 ±1.36 8.85 ±6.21
GraphMAE 31.50 ±11.89 19.93 ±6.60 66.97 ±3.70 37.00 ±3.19 42.68 ±6.91 22.10 ±0.86 8.90 ±6.25
EdgePreGPPT 32.94 ±2.82 26.93 ±1.56 63.45 ±3.63 34.00 ±5.69 36.85 ±3.24 20.42 ±0.81 28.90 ±1.64
EdgePreGprompt 30.55 ±4.59 28.46 ±4.25 63.33 ±4.18 32.57 ±3.31 39.21 ±6.09 20.07 ±0.45 4.01 ±0.69
GCL 51.98 ±3.43 26.68 ±2.95 57.59 ±2.89 30.43 ±8.83 45.20 ±2.20 21.25 ±0.78 14.15 ±5.26
SimGRACE 34.93 ±2.55 29.51 ±2.60 42.87 ±1.57 33.00 ±5.27 41.42 ±6.80 21.58 ±0.84 9.11 ±3.50
ALL-in-one DGI 30.45 ±0.19 21.82 ±4.73 43.60 ±5.24 79.95 ±4.01 62.50 ±12.39 19.83 ±3.14 2.65 ±0.09
GraphMAE 25.02 ±7.58 19.38 ±4.38 46.16 ±15.83 82.43 ±2.84 66.47 ±17.32 15.03 ±3.13 0.24 ±0.00
EdgePreGPPT 30.36 ±13.48 25.83 ±9.32 37.61 ±20.58 88.55 ±3.52 47.98 ±27.78 21.49 ±3.02 1.59 ±1.30
EdgePreGprompt 21.59 ±7.30 18.75 ±1.86 39.36 ±11.45 87.16 ±3.02 46.32 ±29.89 21.23 ±5.64 13.01 ±6.29
GCL 25.63 ±17.68 27.93 ±10.59 42.86 ±7.21 84.75 ±8.86 73.28 ±9.91 20.83 ±8.04 6.70 ±6.01
SimGRACE 15.20 ±9.51 25.19 ±12.95 21.59 ±2.58 78.90 ±2.26 64.63 ±25.59 21.26 ±6.24 4.80 ±5.17
Gprompt DGI 27.81 ±10.28 52.84 ±2.36 37.83 ±5.71 57.24 ±27.27 34.47 ±41.15 26.07 ±4.20 56.59 ±4.05
GraphMAE 66.82 ±3.98 60.07 ±3.98 $6$2.25 74.89 ±35.83 36.73 ±43.95 30.71 ±2.12 76.98 ±3.88
EdgePreGPPT 53.15 ±2.85 36.15 ±5.00 62.18 ±4.88 78.22 ±37.33 37.59 ±45.03 28.76 ±3.95 85.40 ±0.79
EdgePreGprompt 62.63 ±3.26 65.34 ±4.60 67.87 ±2.08 78.01 ±37.38 38.45 ±46.03 21.86 ±2.13 81.85 ±3.16
GCL 69.03 ±3.61 61.27 ±5.37 57.21 ±2.78 65.72 ±31.30 36.38 ±43.47 34.67 ±1.28 57.52 ±4.19
SimGRACE 51.27 ±6.05 66.13 ±1.64 44.78 ±6.52 68.34 ±32.42 39.32 ±47.08 30.06 ±3.36 57.36 ±3.68
GPF DGI 29.57 ±20.89 24.55 ±12.61 52.43 ±6.94 83.21 ±5.19 89.30 ±1.58 35.40 ±8.78 20.72 ±0.26
GraphMAE 35.43 ±1.02 25.12 ±3.01 68.96 ±3.99 96.30 ±5.12 92.44 ±3.04 44.07 ±3.94 45.20 ±2.03
EdgePreGPPT 15.39 ±8.19 18.11 ±2.78 58.67 ±3.21 98.26 ±1.19 89.47 ±1.91 30.99 ±4.68 71.83 ±9.37
EdgePreGprompt 31.58 ±18.16 24.55 ±8.28 66.25 ±8.53 96.30 ±5.12 90.18 ±1.57 31.92 ±5.75 28.60 ±3.11
GCL 28.60 ±11.19 17.69 ±1.35 52.47 ±6.73 69.15 ±21.44 98.42 ±0.36 29.13 ±2.63 23.90 ±0.12
SimGRACE 18.94 ±12.60 22.45 ±3.45 40.35 ±1.71 92.29 ±3.36 91.73 ±3.84 30.28 ±2.64 33.77 ±8.48
GPF-plus DGI 27.23 ±14.61 26.31 ±11.48 47.02 ±14.51 83.86 ±17.24 96.18 ±4.12 36.03 ±7.49 16.86 ±3.30
GraphMAE 63.28 ±4.69 75.73 ±2.19 69.59 ±4.33 99.01 ±1.43 99.12 ±0.95 44.58 ±5.95 47.79 ±1.09
EdgePreGPPT 22.44 ±23.88 13.63 ±5.27 66.43 ±3.28 98.52 ±2.07 96.18 ±4.12 37.15 ±8.48 66.88 ±6.14
EdgePreGprompt 66.22 ±6.20 64.49 ±14.12 68.10 ±4.56 98.64 ±2.14 97.74 ±2.47 41.98 ±4.70 51.25 ±1.31
GCL 47.71 ±22.44 29.16 ±16.66 64.53 ±4.15 70.98 ±19.27 99.12 ±0.95 36.99 ±6.59 25.74 ±1.70
SimGRACE 27.79 ±21.51 28.37 ±22.54 50.25 ±10.04 89.57 ±4.22 95.83 ±3.85 38.23 ±2.34 44.70 ±2.69
Table 18: Node classification performance F1-score (5-shot)
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 51.42 ±8.70 34.20 ±6.12 67.45 ±2.25 34.31 ±4.41 35.02 ±8.10 16.89 ±1.20 14.76 ±1.13
Pretrain + Fine-tuning DGI 47.53 ±9.44 28.71 ±2.76 59.33 ±10.23 34.71 ±5.28 38.80 ±7.04 16.49 ±1.42 4.78 ±2.53
GraphMAE 55.17 ±5.27 29.67 ±6.94 69.58 ±4.63 34.69 ±5.63 34.57 ±9.23 16.92 ±1.76 12.97 ±0.32
EdgePreGPPT 55.50 ±3.62 31.31 ±8.86 70.31 ±4.65 34.86 ±4.83 36.63 ±7.69 19.42 ±1.02 20.19 ±1.27
EdgePredGprompt 49.18 ±8.50 27.04 ±6.83 69.46 ±4.79 34.90 ±5.77 36.38 ±7.04 18.28 ±1.91 18.51 ±1.18
GCL 61.32 ±3.43 35.95 ±3.18 66.16 ±5.43 26.66 ±5.78 33.81 ±3.28 20.44 ±1.45 6.96 ±1.07
SimGRACE 45.33 ±4.88 34.74 ±7.40 62.66 ±6.39 35.86 ±4.86 36.86 ±7.72 20.63 ±1.13 4.74 ±0.48
GPPTPrompt DGI 40.79 ±7.28 39.86 ±7.94 44.67 ±12.25 32.76 ±3.66 39.86 ±4.07 19.28 ±0.85 0.39 ±0.27
GraphMAE 24.55 ±7.97 7.49 ±4.08 66.55 ±3.62 33.86 ±3.88 36.66 ±2.42 19.92 ±0.84 0.42 ±0.29
EdgePreGPPT 32.50 ±2.42 19.95 ±1.98 62.42 ±4.78 30.57 ±4.59 33.93 ±2.76 19.89 ±0.56 18.11 ±1.03
EdgePreGprompt 22.20 ±3.24 25.96 ±4.33 61.22 ±5.46 30.39 ±2.77 35.84 ±3.35 19.22 ±0.33 2.56 ±1.55
GCL 50.60 ±2.51 25.00 ±2.05 57.27 ±2.84 18.78 ±5.00 35.83 ±1.96 20.34 ±0.74 3.03 ±1.03
SimGRACE 31.56 ±2.75 28.27 ±1.68 40.87 ±2.90 27.06 ±3.81 36.11 ±2.81 20.45 ±0.51 1.63 ±0.40
ALL-in-one DGI 6.67 ±0.03 10.68 ±7.84 26.81 ±9.85 60.73 ±6.34 31.59 ±18.97 10.66 ±3.88 0.21 ±0.04
GraphMAE 6.77 ±3.23 8.66 ±4.25 33.97 ±19.42 58.02 ±3.64 49.27 ±9.03 8.31 ±2.80 0.01 ±0.00
EdgePreGPPT 15.19 ±16.25 13.46 ±10.23 31.47 ±24.42 78.97 ±11.06 29.95 ±20.92 10.10 ±3.63 0.53 ±0.63
EdgePreGprompt 4.99 ±1.38 5.28 ±0.47 27.49 ±15.02 68.69 ±3.29 25.44 ±17.47 8.08 ±2.12 0.56 ±0.27
GCL 14.88 ±18.86 14.98 ±12.03 31.71 ±13.56 74.33 ±11.49 49.70 ±21.87 10.10 ±6.01 0.75 ±1.07
SimGRACE 7.21 ±7.14 13.71 ±14.05 13.42 ±3.26 66.26 ±6.68 45.47 ±21.83 11.42 ±5.14 0.52 ±0.70
Gprompt DGI 19.96 ±9.68 47.84 ±2.39 37.26 ±5.45 48.86 ±23.99 29.50 ±36.22 24.02 ±4.65 49.69 ±1.68
GraphMAE 61.57 ±5.38 56.39 ±3.60 65.31 ±2.36 70.18 ±34.72 36.05 ±43.75 30.08 ±1.88 66.82 ±2.68
EdgePreGPPT 51.03 ±3.08 31.15 ±3.86 60.74 ±4.30 73.65 ±36.66 34.07 ±41.83 27.29 ±3.43 79.73 ±1.29
EdgePreGprompt 60.77 ±3.39 63.42 ±5.11 66.46 ±2.30 77.08 ±37.97 34.45 ±41.99 21.03 ±2.04 79.20 ±1.95
GCL 64.22 ±4.08 57.65 ±5.74 56.66 ±2.35 51.82 ±25.62 31.38 ±38.22 32.99 ±1.00 49.82 ±2.81
SimGRACE 45.98 ±5.22 63.84 ±1.37 42.20 ±5.64 62.12 ±30.55 38.47 ±46.70 29.63 ±3.89 53.41 ±2.08
GPF DGI 18.44 ±18.83 13.25 ±11.04 47.23 ±9.28 75.78 ±14.28 77.01 ±10.57 28.58 ±14.86 11.02 ±0.99
GraphMAE 23.17 ±2.53 12.78 ±1.23 67.88 ±4.07 87.75 ±13.24 82.81 ±10.55 43.09 ±4.07 16.76 ±2.76
EdgePreGPPT 3.69 ±1.67 5.09 ±0.66 53.91 ±7.41 94.37 ±8.31 80.86 ±9.83 23.04 ±9.22 55.01 ±8.05
EdgePreGprompt 20.17 ±13.79 12.79 ±6.88 62.47 ±11.25 89.10 ±11.52 79.88 ±10.50 30.44 ±9.87 18.77 ±3.29
GCL 11.74 ±6.74 5.99 ±2.19 46.46 ±14.28 51.01 ±22.81 86.64 ±9.95 22.59 ±5.87 18.74 ±0.78
SimGRACE 10.17 ±12.35 9.91 ±2.56 29.36 ±8.75 71.38 ±2.97 81.14 ±11.49 22.30 ±8.60 12.31 ±4.13
GPF-plus DGI 13.92 ±14.65 12.41 ±13.45 40.73 ±16.83 71.22 ±13.69 84.85 ±10.28 32.05 ±9.40 6.42 ±2.47
GraphMAE 61.55 ±4.90 74.86 ±2.27 68.34 ±4.25 94.48 ±8.79 86.93 ±10.70 44.28 ±6.46 24.14 ±2.16
EdgePreGPPT 15.91 ±26.40 4.26 ±1.11 61.79 ±8.54 93.18 ±10.98 82.10 ±9.29 35.11 ±10.76 56.15 ±4.28
EdgePreGprompt 65.96 ±5.07 57.16 ±18.91 66.11 ±4.85 93.46 ±10.79 85.56 ±9.53 39.84 ±6.02 32.07 ±2.05
GCL 38.84 ±26.82 18.53 ±21.72 63.06 ±3.88 65.13 ±17.26 86.59 ±11.04 33.23 ±7.80 19.46 ±1.47
SimGRACE 19.60 ±23.99 18.27 ±26.81 36.93 ±16.65 69.27 ±9.83 84.63 ±10.23 35.78 ±3.00 24.02 ±2.29
Table 19: Node classification performance AUROC (5-shot)
Training schemes Methods Cora CiteSeer Pubmed Wisconsin Texas Actor Ogbn-arxiv
self-supervise GCN 88.39 ±2.56 78.16 ±2.95 84.93 ±2.77 65.96 ±5.03 66.32 ±8.89 51.44 ±0.62 78.83 ±1.37
Pretrain + Fine-tuning DGI 84.60 ±5.40 74.29 ±3.78 81.26 ±3.74 67.40 ±4.97 66.62 ±8.12 51.78 ±0.48 62.86 ±4.35
GraphMAE 90.30 ±1.62 77.04 ±3.75 86.09 ±2.54 65.32 ±5.26 63.95 ±9.63 51.73 ±0.46 75.65 ±0.78
EdgePreGPPT 86.88 ±1.90 75.17 ±2.72 86.23 ±2.32 67.71 ±3.54 69.36 ±8.52 51.34 ±0.65 83.86 ±1.25
EdgePredGprompt 87.41 ±2.15 73.79 ±5.07 85.66 ±3.26 66.04 ±5.12 66.70 ±8.92 52.40 ±0.91 82.34 ±1.45
GCL 90.49 ±1.49 76.54 ±1.56 83.07 ±3.73 57.71 ±8.14 65.57 ±7.78 52.68 ±0.69 66.70 ±0.95
SimGRACE 85.45 ±2.29 76.95 ±2.79 80.86 ±3.17 67.73 ±4.47 66.11 ±7.69 52.52 ±0.92 60.36 ±0.72
GPPTPrompt DGI 76.38 ±6.02 76.59 ±3.18 63.76 ±10.00 69.27 ±1.49 66.52 ±3.33 51.27 ±1.04 50.09 ±0.25
GraphMAE 84.77 ±3.18 69.69 ±1.62 82.56 ±1.84 66.53 ±4.44 63.43 ±2.99 50.92 ±0.44 63.55 ±2.93
EdgePreGPPT 69.44 ±2.15 60.14 ±1.47 77.28 ±3.58 62.04 ±3.27 61.22 ±2.13 50.85 ±0.37 85.67 ±0.67
EdgePreGprompt 69.60 ±0.83 62.44 ±2.17 81.98 ±4.11 63.83 ±2.92 65.86 ±2.34 50.32 ±0.54 71.01 ±5.38
GCL 82.06 ±2.24 61.54 ±2.42 76.26 ±1.36 55.23 ±2.54 62.83 ±4.62 50.87 ±0.69 58.53 ±1.53
SimGRACE 70.54 ±2.07 64.49 ±1.54 60.67 ±1.80 66.38 ±0.42 62.69 ±1.92 51.00 ±0.54 51.46 ±0.26
ALL-in-one DGI 42.65 ±0.41 57.45 ±7.01 57.38 ±8.87 85.53 ±2.93 68.87 ±6.32 51.24 ±2.32 54.76 ±0.75
GraphMAE 80.75 ±0.64 57.95 ±1.98 81.32 ±2.25 83.91 ±5.14 72.30 ±3.24 53.43 ±0.71 53.46 ±2.41
EdgePreGPPT 62.42 ±11.37 62.25 ±7.23 63.45 ±13.43 95.30 ±4.46 78.68 ±10.06 51.90 ±0.91 54.22 ±1.57
EdgePreGprompt 77.24 ±3.51 64.14 ±1.01 75.97 ±3.33 91.72 ±1.62 80.10 ±10.86 50.40 ±3.16 75.81 ±2.83
GCL 65.99 ±10.83 65.27 ±11.57 69.76 ±4.24 88.87 ±4.34 76.09 ±6.05 54.21 ±2.36 55.77 ±4.78
SimGRACE 59.53 ±3.53 64.55 ±7.97 47.71 ±3.86 86.55 ±2.81 83.01 ±10.84 52.44 ±1.68 53.56 ±7.86
Gprompt DGI 70.76 ±4.67 80.01 ±2.22 54.50 ±2.51 84.53 ±17.37 65.41 ±9.24 54.02 ±2.61 92.66 ±0.73
GraphMAE 72.59 ±7.56 76.04 ±5.23 82.88 ±2.18 89.40 ±19.62 66.34 ±10.38 56.23 ±2.46 96.45 ±0.14
EdgePreGPPT 85.57 ±0.39 68.97 ±4.56 76.57 ±1.08 89.57 ±19.71 66.48 ±10.55 58.85 ±2.98 98.40 ±0.08
EdgePreGprompt 88.49 ±0.88 84.82 ±2.89 83.58 ±1.29 89.71 ±19.78 66.52 ±10.61 52.51 ±2.39 97.79 ±0.15
GCL 88.70 ±2.05 81.00 ±6.61 74.92 ±2.53 83.64 ±16.83 66.33 ±10.37 63.03 ±3.07 94.54 ±0.37
SimGRACE 78.28 ±3.96 89.06 ±0.86 59.22 ±5.76 86.22 ±18.17 66.68 ±10.79 58.88 ±3.13 95.27 ±0.28
GPF DGI 62.20 ±15.24 62.32 ±17.75 67.14 ±11.86 89.86 ±8.25 78.39 ±1.66 65.97 ±7.26 74.59 ±0.97
GraphMAE 80.34 ±2.03 67.33 ±3.16 86.64 ±2.02 97.72 ±3.33 89.25 ±8.10 73.06 ±2.24 74.39 ±0.76
EdgePreGPPT 49.95 ±0.10 52.76 ±6.92 74.91 ±5.59 99.98 ±0.03 83.65 ±7.64 59.31 ±5.01 90.96 ±1.32
EdgePreGprompt 73.56 ±5.48 76.82 ±4.28 85.50 ±1.95 97.15 ±4.98 86.81 ±8.42 60.43 ±6.30 75.36 ±2.61
GCL 79.04 ±5.41 64.25 ±7.12 70.90 ±6.31 73.84 ±18.96 83.62 ±4.75 57.39 ±3.99 75.19 ±0.54
SimGRACE 60.98 ±9.22 64.93 ±3.80 58.21 ±3.99 90.38 ±1.52 90.43 ±8.89 61.68 ±2.02 60.89 ±3.56
GPF-plus DGI 60.30 ±14.39 54.40 ±11.09 73.47 ±9.02 95.86 ±3.95 89.08 ±8.43 68.06 ±2.70 67.37 ±3.03
GraphMAE 89.99 ±1.39 92.09 ±1.11 85.69 ±3.12 99.93 ±0.14 87.29 ±7.92 73.10 ±2.68 81.99 ±0.56
EdgePreGPPT 65.26 ±14.38 55.10 ±2.95 82.94 ±2.55 99.96 ±0.05 86.89 ±8.72 70.17 ±4.50 93.73 ±1.05
EdgePreGprompt 91.09 ±1.03 82.22 ±12.12 83.84 ±3.91 99.96 ±0.08 83.88 ±7.87 72.25 ±1.48 81.54 ±1.05
GCL 84.11 ±7.92 61.55 ±15.26 81.38 ±3.75 91.95 ±9.63 86.24 ±7.72 69.74 ±1.87 77.42 ±2.14
SimGRACE 66.55 ±14.49 63.11 ±14.51 65.07 ±10.97 88.17 ±7.54 89.58 ±9.03 68.81 ±1.60 69.60 ±2.37
Table 20: Graph classification performance Accuracy (%) (1-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 57.30±0.98 47.23±0.61 56.36±7.97 65.20±6.70 20.58±2.00 27.08±11.94 25.80±6.53 55.33±6.22
Pretrain + Fine-tuning DGI 57.32±0.90 42.22±0.73 60.00±4.48 64.13±7.90 17.83±1.88 29.44±9.68 26.48±7.61 57.15±4.32
GraphMAE 57.70±1.13 48.10±0.23 62.40±1.94 65.20±5.00 22.21±2.79 28.47±14.72 25.80±6.53 53.59±6.93
EdgePreGPPT 57.20±0.85 47.14±0.55 58.27±10.66 64.27±4.73 19.79±2.17 27.83±13.44 34.69±8.50 52.82±9.38
EdgePreGprompt 57.35±0.92 47.20±0.53 61.84±2.59 62.67±2.67 19.75±2.33 27.13±12.05 29.44±11.20 56.16±5.10
GCL 57.75±1.02 39.62±0.63 63.44±3.64 65.07±8.38 21.21±0.87 53.14±21.32 29.07±7.00 55.50±5.83
SimGRACE 57.33±0.96 46.89±0.42 60.07±3.21 65.47±5.89 19.71±1.76 76.19±5.41 28.46±6.49 53.23±9.71
GPPTPrompt DGI 49.07±10.36 39.34±9.11 60.81±1.55 51.33±15.87 20.29±1.40 78.23±1.38 44.07±22.42 53.65±10.00
GraphMAE 50.15±0.75 29.46±13.65 60.72±1.70 44.80±15.52 20.37±1.96 68.63±20.51 55.99±19.28 57.69±6.89
EdgePreGPPT 49.38±10.29 36.47±7.88 60.92±2.47 42.80±12.98 20.87±2.42 73.99±9.79 49.81±20.17 53.69±6.88
EdgePreGprompt 50.15±0.75 40.22±9.56 57.03±4.55 37.87±10.43 22.08±3.42 72.28±13.22 50.06±18.97 55.33±8.51
GCL 45.70±8.20 47.18±5.93 59.24±1.01 60.40±15.43 21.29±3.79 68.36±21.05 59.32±11.22 56.26±8.20
SimGRACE 46.03±10.29 41.11±8.47 55.42±8.81 52.67±17.12 20.83±3.47 62.31±19.42 59.20±15.06 55.88±7.81
ALL-in-one DGI 60.07±4.81 39.56±5.00 62.58±7.07 75.73±7.75 23.96±1.45 50.72±9.93 79.20±1.65 55.97±6.52
GraphMAE 52.62±3.04 40.82±14.63 66.49±6.26 72.27±8.87 23.21±1.72 56.68±7.38 75.37±5.99 58.77±1.05
EdgePreGPPT 59.12±0.77 42.74±4.65 65.71±5.49 79.87±5.34 20.92±2.04 60.27±16.97 64.20±19.74 56.24±2.46
EdgePreGprompt 53.78±2.82 42.87±6.19 61.82±7.53 75.47±5.00 21.88±0.56 49.06±5.53 66.60±17.36 57.60±4.37
GCL 58.75±0.80 51.66±0.26 64.36±7.30 70.53±8.58 19.46±2.85 52.55±13.51 50.93±16.83 59.72±1.52
SimGRACE 58.83±0.85 47.60±0.39 61.17±1.73 65.47±6.91 22.50±1.56 76.14±5.51 69.88±18.06 58.26±1.18
Gprompt DGI 50.47±10.10 47.29±7.78 56.61±7.93 63.33±14.36 20.50±1.79 45.52±16.98 55.43±13.69 56.18±6.13
GraphMAE 54.75±12.43 36.39±7.72 57.66±12.56 68.80±4.76 19.54±1.99 43.91±6.64 47.16±4.72 55.22±6.40
EdgePreGPPT 51.18±11.11 46.70±5.74 59.17±11.26 52.13±5.80 19.71±4.46 50.08±8.00 45.06±15.93 51.04±4.82
EdgePreGprompt 51.57±11.87 40.53±12.02 55.55±8.17 73.60±4.76 19.67±3.08 54.64±9.94 51.36±15.55 57.20±5.54
GCL 50.50±10.42 45.54±9.05 55.51±10.73 56.00±13.79 19.83±2.19 44.40±5.74 46.42±20.67 52.65±9.17
SimGRACE 50.40±10.54 48.25±13.64 57.53±11.05 64.67±7.92 22.29±3.50 47.02±5.59 52.90±11.76 57.81±2.68
GPF DGI 52.85±8.91 42.75±7.10 59.17±3.63 63.07±7.22 22.00±1.25 27.94±13.65 70.56±15.46 59.36±1.18
GraphMAE 49.27±7.77 37.23±17.95 58.65±8.49 65.73±6.91 20.71±1.92 40.43±10.43 67.84±20.28 55.80±6.34
EdgePreGPPT 59.35±1.02 37.53±5.19 62.54±2.55 66.40±5.93 22.04±1.48 27.40±12.58 24.75±8.01 43.86±5.37
EdgePreGprompt 59.65±5.06 41.44±0.52 61.82±2.61 68.40±5.09 22.17±1.48 65.79±17.72 60.49±24.20 55.56±5.07
GCL 57.73±0.79 47.42±1.22 63.91±3.26 59.20±7.01 21.13±2.11 37.05±9.39 71.67±14.71 58.43±1.17
SimGRACE 58.30±0.77 41.04±0.24 63.35±3.69 66.93±5.05 21.79±2.40 33.99±11.05 27.47±5.92 58.03±1.63
GPF-plus DGI 57.87±6.19 43.98±7.17 61.26±3.06 62.53±6.86 18.71±1.19 26.70±11.19 67.90±14.60 48.79±9.14
GraphMAE 55.00±5.81 40.32±9.49 62.49±2.05 62.13±8.71 22.92±1.64 33.78±11.61 68.40±20.02 56.31±4.93
EdgePreGPPT 56.65±4.08 41.41±0.73 63.06±2.55 65.07±4.41 21.50±2.37 27.83±11.54 29.57±7.69 57.62±2.42
EdgePreGprompt 55.35±3.91 40.08±2.95 61.33±2.81 65.20±6.04 19.42±1.88 28.85±15.48 57.22±18.17 55.90±5.56
GCL 57.10±1.34 46.89±0.24 59.75±7.95 64.00±7.89 18.79±1.46 25.90±9.58 71.17±14.92 57.56±2.54
SimGRACE 57.93±1.62 47.24±0.29 62.92±2.78 61.33±3.84 20.29±1.62 27.08±7.86 28.46±6.49 57.11±3.42
Table 21: Graph classification performance F1-score (1-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 54.62±1.12 41.10±0.39 46.69±10.82 63.47±6.36 15.25±3.96 22.78±10.69 23.71±8.23 44.74±4.23
Pretrain + Fine-tuning DGI 54.60±1.00 38.53±0.34 54.82±3.34 61.97±7.76 10.76±4.28 27.09±11.48 24.34±9.21 46.15±5.41
GraphMAE 55.20±1.24 41.71±0.17 52.05±7.26 63.41±4.44 19.17±3.42 23.63±12.40 23.71±8.23 46.25±7.84
EdgePreGPPT 54.39±0.95 41.10±0.37 55.82±10.61 60.94±3.46 12.89±3.54 23.08±11.30 33.12±7.45 36.31±5.78
EdgePreGprompt 54.62±1.07 41.14±0.41 59.73±1.34 59.05±1.33 13.72±4.13 22.91±10.95 27.09±12.51 45.49±4.58
GCL 55.24±1.07 36.27±0.63 56.25±7.55 63.37±8.64 16.78±1.91 39.11±4.29 27.67±8.70 48.68±6.42
SimGRACE 54.69±1.09 40.92±0.37 52.67±7.14 63.70±5.32 14.15±2.49 45.06±1.93 27.05±8.20 37.84±7.14
GPPTPrompt DGI 41.17±12.71 27.05±13.23 46.05±10.61 41.76±15.84 17.26±2.39 44.68±1.17 33.93±14.06 43.61±5.55
GraphMAE 33.40±0.33 14.61±5.30 46.64±11.32 39.44±18.49 17.61±2.22 40.48±7.24 44.38±9.91 50.34±5.80
EdgePreGPPT 44.16±6.70 21.35±9.96 47.07±11.95 38.03±16.75 17.02±2.90 43.67±0.88 40.73±9.69 51.50±6.54
EdgePreGprompt 33.40±0.33 18.92±3.19 43.34±8.00 31.42±13.60 19.87±2.99 43.73±0.75 41.78±10.08 45.12±7.86
GCL 39.08±10.25 42.87±7.70 41.15±7.80 53.15±16.82 19.62±3.92 40.11±7.97 45.58±3.38 50.02±8.57
SimGRACE 43.18±8.95 33.88±13.05 40.87±7.11 46.54±18.13 18.87±3.37 41.86±9.35 49.40±8.41 46.82±6.89
ALL-in-one DGI 56.82±6.07 35.40±5.55 60.66±6.94 74.53±7.04 14.48±3.58 44.46±4.45 62.11±7.06 48.28±7.29
GraphMAE 45.83±5.38 18.76±5.47 64.27±4.78 69.78±8.83 19.66±3.11 49.40±3.96 53.85±3.84 56.70±1.89
EdgePreGPPT 57.29±0.74 37.07±5.56 64.68±5.35 78.57±4.92 12.95±3.18 49.62±10.42 51.66±12.74 55.10±1.49
EdgePreGprompt 48.44±4.51 34.64±5.55 60.04±8.57 73.31±5.46 12.50±3.12 45.57±5.70 44.06±5.14 48.13±4.31
GCL 56.83±0.76 47.78±0.10 62.99±7.19 64.90±11.24 12.01±5.16 46.65±6.50 41.89±8.44 43.55±8.21
SimGRACE 56.88±0.80 41.64±0.13 53.18±7.57 58.57±11.26 12.23±2.42 45.03±1.86 46.16±7.12 39.55±5.05
Gprompt DGI 48.68±9.78 42.80±9.19 55.95±7.78 61.15±13.98 18.68±2.94 38.30±12.89 44.61±5.71 49.81±1.61
GraphMAE 52.10±13.61 17.64±2.56 55.24±12.01 64.58±3.26 18.36±2.20 42.68±5.98 43.38±3.73 50.47±3.41
EdgePreGPPT 49.33±10.58 43.20±8.14 58.30±10.88 50.70±6.00 18.20±5.07 44.54±3.28 39.06±9.23 50.78±5.00
EdgePreGprompt 50.43±11.93 36.62±12.55 54.29±7.32 71.38±3.64 17.17±4.25 46.26±5.14 43.73±9.27 48.18±4.55
GCL 48.91±10.12 40.78±10.09 53.98±9.93 53.39±14.36 18.26±2.77 42.26±4.15 38.58±11.82 50.85±8.14
SimGRACE 48.78±10.20 43.35±10.75 55.51±10.10 60.58±6.08 19.52±3.36 44.68±4.01 44.81±6.73 52.80±3.60
GPF DGI 50.50±7.88 34.56±6.02 49.27±10.07 62.02±6.87 15.08±1.44 23.42±11.99 44.77±3.37 39.53±5.01
GraphMAE 42.98±8.29 17.19±6.84 52.62±9.40 59.14±10.41 13.10±3.28 37.75±10.47 41.84±7.65 48.52±7.11
EdgePreGPPT 55.67±0.84 34.09±5.32 57.01±5.79 58.18±3.05 16.57±1.64 22.68±10.50 21.77±9.16 34.22±10.09
EdgePreGprompt 56.22±6.17 38.14±0.44 56.91±6.21 63.90±4.05 17.34±2.45 43.08±4.88 39.86±11.54 47.44±4.83
GCL 55.23±0.77 38.04±0.46 56.08±7.40 57.99±6.96 15.97±3.75 35.89±9.97 48.83±5.30 40.86±4.89
SimGRACE 56.19±0.68 37.69±0.21 55.50±9.14 58.38±2.44 14.39±3.45 31.82±12.08 26.02±7.56 39.13±4.20
GPF-plus DGI 53.13±10.49 37.59±1.42 54.74±7.01 61.19±6.31 13.03±1.21 22.69±10.52 46.57±4.62 33.21±4.87
GraphMAE 48.23±10.59 18.94±3.14 52.88±6.59 61.01±8.89 18.39±2.76 30.90±11.56 44.87±9.19 46.24±4.86
EdgePreGPPT 50.88±7.65 37.29±2.27 57.58±7.28 62.03±2.92 17.40±2.32 24.17±10.38 28.06±9.23 40.06±6.06
EdgePreGprompt 50.07±6.94 37.43±1.84 54.79±2.74 63.20±5.31 14.44±1.44 24.60±14.35 42.77±1.25 45.59±4.64
GCL 54.24±1.69 38.53±0.20 57.54±6.94 62.31±7.93 13.66±2.70 22.64±10.42 48.71±5.51 39.37±4.68
SimGRACE 55.55±2.03 41.24±0.31 54.80±6.88 59.38±2.64 14.74±2.94 24.79±9.43 27.05±8.20 39.51±4.96
Table 22: Graph classification performance AUROC (1-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 67.05±1.01 54.23±0.34 57.88±1.72 71.68±1.25 53.49±1.11 48.39±1.89 51.13±1.38 49.60±2.94
Pretrain + Fine-tuning DGI 67.06±1.00 74.13±1.27 56.87±2.74 71.45±1.55 51.76±2.18 52.35±7.49 53.79±1.12 49.32±3.05
GraphMAE 66.91±1.18 55.16±0.20 59.87±0.78 69.81±1.39 53.43±0.98 49.46±1.08 51.01±0.87 50.66±4.90
EdgePreGPPT 67.50±1.06 54.41±0.31 58.29±3.43 72.42±1.64 53.62±1.61 46.88±1.16 48.35±3.28 52.27±1.89
EdgePreGprompt 66.92±0.96 54.36±0.30 58.24±1.27 71.58±0.75 53.30±1.75 49.62±0.75 55.48±2.09 49.97±3.16
GCL 67.11±1.09 68.95±1.48 59.68±1.80 71.07±1.87 54.35±0.92 51.84±2.24 53.34±2.75 53.14±5.70
SimGRACE 66.95±0.99 53.79±0.47 59.80±1.06 71.49±1.82 54.20±1.36 49.38±1.24 54.01±2.14 49.20±0.84
GPPTPrompt DGI 48.58±10.92 58.79±4.61 71.14±3.06 51.11±31.50 53.09±1.46 53.40±0.84 48.98±7.81 53.82±11.83
GraphMAE 50.27±0.74 51.39±4.26 70.25±3.45 36.15±27.67 52.89±1.24 53.91±1.86 54.53±11.66 58.36±8.51
EdgePreGPPT 48.81±11.41 56.39±0.58 70.36±3.79 33.54±24.51 52.98±1.23 50.31±5.34 51.48±6.83 54.80±8.30
EdgePreGprompt 50.27±0.74 54.48±1.95 65.03±10.25 24.49±23.32 53.83±2.13 51.51±2.95 53.38±9.22 56.26±10.88
GCL 45.42±8.63 67.97±6.02 69.82±6.05 65.58±22.89 53.36±2.88 54.61±3.27 51.68±3.28 56.31±9.89
SimGRACE 45.58±10.97 61.78±7.34 62.48±15.87 48.86±29.41 53.23±2.42 55.27±7.83 58.65±6.19 56.15±9.67
ALL-in-one DGI 69.12±1.04 66.27±11.54 75.07±0.76 90.86±1.09 57.42±1.15 50.21±1.87 57.93±12.50 60.56±0.91
GraphMAE 65.65±1.22 50.00±0.00 73.97±0.80 86.98±0.90 54.66±0.90 57.09±6.45 56.21±11.12 59.65±2.74
EdgePreGPPT 65.44±0.92 64.21±9.84 75.51±0.71 91.34±0.84 55.67±1.65 61.50±1.13 59.90±1.41 57.44±0.59
EdgePreGprompt 68.48±1.11 51.35±0.30 73.73±0.94 86.05±1.03 55.15±0.59 57.02±10.92 54.72±1.48 56.95±4.42
GCL 65.20±1.06 63.82±0.26 77.76±0.74 85.11±0.98 53.58±1.38 66.43±2.88 49.56±2.28 55.04±1.51
SimGRACE 66.33±0.99 53.09±0.21 57.32±1.26 74.91±1.00 54.75±0.64 45.81±0.74 55.96±0.33 50.64±1.32
Gprompt DGI 53.74±19.12 73.04±8.22 60.36±9.11 69.64±13.88 55.68±2.50 49.63±0.56 49.08±3.10 52.16±4.08
GraphMAE 51.48±9.49 44.86±0.14 55.65±7.44 70.27±4.19 55.28±3.35 56.97±3.76 51.24±5.39 52.47±6.09
EdgePreGPPT 55.21±19.21 69.85±10.59 60.31±14.12 59.50±10.08 53.75±2.67 53.86±6.62 49.52±5.86 52.73±6.44
EdgePreGprompt 53.58±19.20 73.09±5.29 57.72±7.86 79.17±2.18 55.53±3.27 48.74±2.54 50.81±7.50 54.96±5.21
GCL 53.54±18.02 66.59±11.21 57.80±12.67 71.02±3.35 55.06±3.09 54.08±4.37 47.73±6.12 52.92±6.67
SimGRACE 53.21±18.90 66.59±8.84 59.42±13.20 66.20±4.91 55.23±3.64 53.95±3.70 49.34±3.63 52.15±6.51
GPF DGI 55.96±12.77 50.91±1.05 59.34±0.55 71.40±1.24 53.13±1.13 48.97±0.73 47.84±2.93 49.47±0.63
GraphMAE 61.68±14.66 50.00±0.00 59.21±1.29 73.95±2.06 51.73±1.39 52.21±1.49 46.87±1.80 49.36±4.14
EdgePreGPPT 69.73±3.23 64.68±11.19 61.63±2.41 68.41±1.95 54.48±1.10 48.80±2.39 51.30±2.60 49.08±3.86
EdgePreGprompt 69.06±1.34 71.37±1.86 58.36±1.20 74.51±1.18 54.01±0.86 50.42±0.62 58.34±0.87 52.14±2.73
GCL 67.60±1.02 54.30±0.39 59.15±1.73 68.50±2.13 53.76±1.38 59.21±1.87 56.57±1.65 49.30±1.63
SimGRACE 66.95±1.01 73.55±0.37 60.00±1.43 68.25±1.28 55.25±1.53 55.80±0.69 52.30±1.91 46.17±0.78
GPF-plus DGI 68.77±1.50 72.06±4.49 58.37±2.32 70.12±1.06 53.83±0.61 48.71±0.50 51.49±1.31 48.76±1.56
GraphMAE 68.98±1.49 50.00±0.00 59.58±1.13 72.03±2.16 52.68±1.28 47.21±1.67 53.26±1.91 49.42±2.18
EdgePreGPPT 69.57±2.53 69.92±1.65 63.58±3.31 72.02±0.86 55.66±1.62 45.11±0.89 55.06±4.29 44.29±3.43
EdgePreGprompt 68.39±1.07 69.85±1.36 57.57±1.65 73.37±1.78 52.19±1.18 51.14±2.27 56.22±5.10 49.74±1.32
GCL 67.01±1.22 54.03±0.40 59.78±1.47 71.25±1.88 52.59±2.06 51.25±2.24 56.70±1.44 48.94±2.26
SimGRACE 67.19±1.54 53.04±0.17 59.66±1.32 68.33±0.76 52.47±1.76 51.79±1.73 53.39±2.81 49.75±1.52
Table 23: Graph classification performance Accuracy (%) (3-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 53.33±6.61 50.77±2.44 61.33±2.89 59.47±8.34 15.96±1.64 65.15±18.61 52.35±8.12 59.77±1.10
Pretrain + Fine-tuning DGI 53.33±6.61 56.10±3.46 61.33±2.75 59.87±8.78 21.71±0.81 51.96±13.00 52.22±10.64 59.70±0.98
GraphMAE 53.33±6.61 49.11±16.81 61.19±1.57 44.00±13.56 15.04±1.86 60.11±18.73 38.52±17.15 57.90±2.69
EdgePreGPPT 63.43±2.65 47.40±15.87 62.72±2.39 43.47±13.44 21.96±2.45 49.81±9.44 43.70±17.89 58.94±0.66
EdgePreGprompt 53.33±6.61 54.67±1.34 60.67±2.32 59.20±7.05 17.58±1.45 35.82±14.38 29.07±9.83 59.45±9.10
GCL 62.22±1.38 55.27±2.61 62.07±2.39 54.80±5.97 22.00±1.71 31.90±8.27 51.23±11.67 55.54±6.26
SimGRACE 66.10±0.70 55.38±3.58 60.09±0.63 54.00±7.03 22.71±0.86 69.97±13.89 36.67±1.80 58.17±2.79
GPPTPrompt DGI 50.33±0.92 38.40±8.13 60.36±8.99 64.13±18.31 17.67±2.05 56.84±28.02 67.53±23.14 56.11±7.51
GraphMAE 50.23±0.95 36.37±7.89 56.94±6.67 47.87±17.55 17.63±1.97 69.38±19.31 70.49±17.22 56.50±8.48
EdgePreGPPT 59.48±5.42 38.45±9.13 64.74±1.99 52.00±16.93 19.12±2.43 69.87±18.34 58.95±12.79 52.14±6.23
EdgePreGprompt 51.85±2.10 36.11±7.73 60.76±1.52 62.00±18.69 18.71±5.10 52.17±15.58 77.96±2.43 57.94±7.47
GCL 50.43±11.80 50.88±6.31 60.31±0.57 48.93±18.86 17.25±1.19 71.90±14.28 78.46±1.57 56.28±7.21
SimGRACE 50.12±12.86 41.87±8.73 55.93±6.16 57.87±20.52 15.62±2.32 58.28±20.18 67.47±19.05 59.00±6.34
ALL-in-one DGI 64.28±0.75 52.63±8.14 69.84±6.02 82.40±3.57 22.87±0.93 52.17±12.81 59.81±15.62 54.95±6.52
GraphMAE 63.88±0.73 52.09±0.33 65.69±3.31 82.00±4.02 21.04±2.51 53.83±7.02 61.98±11.32 56.56±4.54
EdgePreGPPT 63.80±1.07 55.73±3.59 65.62±7.36 85.60±2.22 22.17±2.17 50.19±10.89 54.26±13.30 52.19±6.17
EdgePreGprompt 63.90±1.57 51.69±8.39 61.30±2.47 73.60±4.23 23.25±1.11 60.21±8.86 54.44±17.27 58.96±5.93
GCL 65.67±0.58 57.12±1.99 65.57±2.24 58.93±10.23 23.96±0.62 52.17±14.65 58.64±4.86 52.65±5.87
SimGRACE 64.20±1.29 55.48±3.48 62.36±1.86 72.40±4.01 22.58±1.18 66.06±18.23 61.30±16.21 53.23±6.95
Gprompt DGI 58.95±9.88 55.27±8.86 62.43±4.09 54.93±17.15 20.50±2.36 50.29±7.71 49.69±6.79 53.84±5.72
GraphMAE 59.17±10.00 36.11±7.73 61.98±4.45 64.40±16.46 21.42±2.71 44.83±9.16 48.15±8.09 52.89±5.27
EdgePreGPPT 64.35±1.21 53.20±7.90 64.94±2.92 53.60±14.41 18.50±3.69 45.47±7.03 54.63±2.95 53.61±2.31
EdgePreGprompt 59.30±10.17 54.95±9.47 62.02±3.15 66.53±14.84 21.42±3.01 50.56±9.27 49.88±12.32 52.55±6.16
GCL 59.85±9.50 52.52±9.71 58.49±9.20 52.40±20.58 21.42±0.77 48.15±9.42 54.26±9.10 55.61±3.21
SimGRACE 60.00±9.95 53.45±7.18 60.27±4.44 56.40±13.37 22.08±3.57 51.53±13.08 43.21±8.84 55.99±7.53
GPF DGI 63.53±2.47 49.84±7.48 61.39±2.63 48.67±15.53 16.63±3.49 65.31±19.45 61.79±21.19 59.07±0.65
GraphMAE 62.80±2.90 37.01±13.81 62.72±3.07 55.87±12.48 18.29±2.39 53.51±13.09 51.91±8.73 57.15±5.68
EdgePreGPPT 65.25±2.65 51.91±8.13 63.35±2.45 74.27±1.55 19.92±2.19 44.50±4.46 54.63±10.59 51.59±5.62
EdgePreGprompt 64.05±1.03 50.37±7.25 62.49±2.18 55.60±13.42 23.08±3.11 61.72±11.67 74.38±11.62 56.37±6.77
GCL 63.25±2.36 53.87±3.44 62.90±2.52 54.00±12.02 22.38±1.93 49.33±11.40 50.19±4.33 52.34±6.89
SimGRACE 65.97±0.69 53.23±4.59 60.92±1.65 50.13±13.88 23.87±3.45 62.31±8.87 25.62±8.25 57.54±4.65
GPF-plus DGI 62.45±2.52 52.14±7.67 62.16±2.14 75.20±3.64 21.92±0.74 65.25±18.07 60.86±16.47 59.43±0.52
GraphMAE 61.97±2.88 36.87±13.90 63.55±1.85 59.33±7.66 17.08±1.68 48.20±19.15 40.99±10.64 57.18±5.31
EdgePreGPPT 64.38±2.30 54.63±7.14 62.99±1.94 72.40±2.00 23.62±2.56 49.44±11.92 36.85±19.33 58.85±1.33
EdgePreGprompt 64.00±3.54 50.77±9.01 60.38±2.47 50.27±18.19 24.46±2.27 52.87±12.00 50.06±16.36 58.54±2.38
GCL 63.25±2.63 56.50±3.71 60.56±1.94 74.27±4.59 19.00±2.42 51.58±11.78 71.67±14.87 53.99±6.38
SimGRACE 63.55±2.25 52.72±6.39 60.07±0.97 50.67±17.37 22.17±2.30 63.86±10.00 25.62±8.25 59.51±0.62
Table 24: Graph classification performance F1-score (3-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 39.88±13.07 40.73±1.00 51.01±8.18 57.79±8.44 11.97±0.92 42.57±5.89 47.71±5.27 41.05±7.70
Pretrain + Fine-tuning DGI 39.88±13.07 56.10±3.90 50.89±8.01 58.19±8.98 14.68±2.46 47.30±8.88 46.39±5.60 41.31±8.21
GraphMAE 39.88±13.08 47.71±19.43 51.90±7.98 37.98±16.20 10.19±2.54 43.74±6.72 33.85±14.41 40.24±6.08
EdgePreGPPT 60.66±4.52 35.28±13.12 54.53±4.29 37.20±15.67 15.47±3.45 45.43±4.85 39.01±11.37 40.78±7.17
EdgePreGprompt 39.88±13.07 35.69±7.57 49.31±7.52 57.28±6.35 12.99±2.98 33.02±14.02 26.66±11.36 48.50±10.32
GCL 60.72±1.89 54.45±4.00 52.75±6.77 53.13±5.12 17.00±3.88 30.17±7.80 45.33±7.28 51.94±4.43
SimGRACE 65.91±0.61 54.77±3.76 48.85±6.10 52.10±6.02 16.13±2.28 45.75±2.10 36.07±1.29 50.54±6.18
GPPTPrompt DGI 33.47±0.41 25.72±11.59 56.81±8.37 56.46±18.74 7.26±4.98 33.97±12.73 38.84±10.65 52.76±6.80
GraphMAE 33.43±0.42 17.63±2.61 44.10±8.25 41.20±19.70 7.70±5.86 41.47±5.35 42.54±3.25 53.64±7.98
EdgePreGPPT 54.15±12.01 28.60±16.18 61.46±1.10 49.73±17.65 12.26±6.37 41.87±4.54 49.38±6.09 50.03±5.25
EdgePreGprompt 36.01±4.61 17.54±2.57 51.33±7.17 61.15±18.28 11.46±8.27 41.89±5.55 46.42±4.52 53.54±7.77
GCL 46.73±14.15 49.62±6.77 41.87±8.55 41.46±20.04 9.53±5.93 43.94±0.52 44.66±1.04 53.74±6.42
SimGRACE 47.67±15.00 38.51±10.60 54.02±4.98 55.79±19.08 6.44±2.25 42.49±8.52 42.36±6.71 56.28±5.60
ALL-in-one DGI 63.99±0.86 50.43±9.08 68.85±5.72 81.20±3.18 14.63±1.31 43.21±7.33 53.56±10.74 43.35±5.95
GraphMAE 62.94±1.54 22.83±0.10 62.48±1.62 79.84±3.94 15.61±1.28 46.34±2.82 55.20±8.12 54.23±4.13
EdgePreGPPT 62.99±2.30 55.56±4.18 64.50±6.98 83.88±2.07 17.11±3.24 42.10±6.16 44.01±4.11 49.69±6.19
EdgePreGprompt 62.88±3.31 49.83±10.02 52.05±8.59 71.06±3.19 17.70±1.81 50.66±4.25 47.08±8.47 55.21±4.94
GCL 65.60±0.57 56.55±1.88 62.62±3.77 57.24±11.36 22.37±1.31 45.27±9.63 49.39±2.68 50.28±4.51
SimGRACE 63.50±2.41 55.34±3.89 60.36±1.11 69.33±2.34 18.20±1.04 41.93±5.18 49.00±9.10 49.06±6.95
Gprompt DGI 58.08±10.23 55.32±8.74 58.05±3.97 52.54±17.09 18.81±1.44 44.13±4.49 45.80±7.78 50.97±5.00
GraphMAE 58.01±10.19 17.54±2.57 56.27±5.10 60.26±15.18 19.39±1.66 42.42±7.00 43.14±4.47 49.67±2.97
EdgePreGPPT 63.89±1.11 53.51±7.87 61.58±3.09 49.72±12.38 16.84±3.73 42.05±5.59 47.32±2.37 52.83±1.88
EdgePreGprompt 58.51±10.62 55.39±8.86 56.27±2.08 64.64±13.86 18.56±3.11 43.87±4.26 44.57±9.37 49.71±7.25
GCL 58.83±10.22 52.04±10.22 57.08±8.81 51.26±19.96 19.60±1.34 44.29±5.52 47.69±4.92 54.82±2.66
SimGRACE 59.21±10.45 52.82±8.34 55.53±7.37 52.05±11.07 21.63±3.19 43.90±7.83 40.90±6.25 53.39±6.72
GPF DGI 62.80±2.95 47.97±9.36 51.68±7.67 44.82±12.44 11.70±4.44 41.60±6.58 40.70±4.95 40.29±5.61
GraphMAE 61.83±3.65 17.49±5.13 55.21±7.59 52.99±14.80 15.77±1.71 45.77±7.77 47.24±5.79 54.79±4.60
EdgePreGPPT 63.82±3.68 50.28±10.09 56.68±3.32 68.47±1.51 18.77±2.43 42.92±3.39 48.61±4.66 49.22±6.62
EdgePreGprompt 63.59±0.99 48.60±9.17 53.75±5.47 51.63±14.57 16.62±1.75 53.47±7.31 48.42±3.50 53.86±7.88
GCL 62.47±2.84 53.42±3.86 55.24±8.26 50.93±14.25 15.40±2.55 45.37±9.89 47.00±2.24 48.53±5.59
SimGRACE 65.85±0.61 52.54±5.59 49.67±6.21 47.58±13.28 18.36±4.95 51.59±4.59 22.84±9.35 55.12±4.27
GPF-plus DGI 60.90±3.22 50.49±9.05 53.28±5.40 73.17±3.16 17.51±1.75 43.17±5.33 46.21±5.60 40.98±7.58
GraphMAE 60.30±3.23 17.43±5.19 59.82±4.60 57.10±7.36 12.55±0.72 37.20±9.82 38.17±8.53 52.72±3.75
EdgePreGPPT 63.35±3.33 53.82±8.44 59.39±1.65 65.96±1.23 18.73±2.85 46.05±10.35 31.44±17.20 41.26±6.70
EdgePreGprompt 63.65±3.51 48.46±10.27 49.63±6.76 48.24±17.45 19.80±4.12 47.82±9.03 43.31±12.09 44.08±7.81
GCL 62.49±3.11 56.38±3.66 49.25±7.67 72.21±3.78 15.66±2.56 47.78±9.88 43.52±1.30 43.72±6.72
SimGRACE 62.91±2.78 51.40±8.33 48.38±6.37 48.42±16.33 16.25±3.30 50.53±3.93 22.82±9.32 40.63±6.86
Table 25: Graph classification performance AUROC (3-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 53.85±7.70 54.29±0.68 58.83±2.09 73.36±0.52 51.34±1.42 53.03±0.66 55.36±2.82 50.97±1.94
Pretrain + Fine-tuning DGI 53.85±7.71 79.16±1.29 59.21±1.87 72.49±0.69 55.06±0.72 59.34±3.68 55.52±3.89 50.91±1.85
GraphMAE 53.86±7.71 72.58±11.44 59.12±1.77 73.70±2.39 51.64±1.15 56.52±2.87 57.30±2.71 50.87±1.75
EdgePreGPPT 68.90±1.00 53.70±2.76 59.61±1.05 63.78±11.40 56.63±1.88 58.62±4.95 53.24±1.06 50.87±1.74
EdgePreGprompt 53.88±7.76 54.66±1.29 57.43±1.26 72.67±0.71 52.81±1.23 56.43±2.54 52.85±3.05 55.71±3.36
GCL 66.91±0.96 77.24±5.27 59.63±1.25 71.94±1.69 55.22±0.62 55.65±1.39 52.36±4.70 56.27±3.63
SimGRACE 69.07±0.74 78.72±0.81 54.56±4.50 71.53±1.53 57.45±1.25 54.90±1.47 53.79±5.86 58.14±3.45
GPPTPrompt DGI 50.70±0.98 59.06±5.42 63.21±13.02 69.64±24.34 51.97±1.01 51.15±3.38 53.46±4.77 58.16±7.86
GraphMAE 50.02±1.20 55.37±1.65 64.79±12.01 36.64±30.88 51.88±0.85 51.90±3.51 57.33±2.22 57.48±8.59
EdgePreGPPT 60.13±5.46 59.32±8.93 68.26±2.20 53.05±23.65 52.97±1.80 54.30±1.55 53.75±1.34 54.71±6.13
EdgePreGprompt 51.09±3.03 55.18±1.81 68.92±2.46 66.15±23.60 51.81±3.18 50.12±2.64 56.67±1.07 59.80±7.91
GCL 50.36±13.14 70.37±8.98 71.44±3.56 41.39±32.92 51.66±0.70 54.79±2.46 55.63±1.51 58.02±8.00
SimGRACE 50.30±13.72 63.01±7.00 57.69±9.30 59.52±26.16 49.66±2.26 53.90±4.39 53.94±3.10 60.31±6.86
ALL-in-one DGI 68.29±0.65 81.02±0.32 78.78±0.89 91.14±1.01 55.50±0.41 50.11±4.43 64.10±1.82 47.72±1.00
GraphMAE 67.69±0.61 50.00±0.00 71.38±4.27 90.04±1.80 54.06±0.38 51.31±6.59 62.34±1.48 59.05±3.75
EdgePreGPPT 67.39±0.61 80.54±1.13 71.08±8.07 92.40±0.68 55.76±1.18 47.76±1.24 48.35±3.93 54.96±0.53
EdgePreGprompt 68.01±0.66 77.37±3.38 59.05±4.06 85.58±1.01 57.14±0.79 56.46±1.24 64.23±1.73 60.92±1.89
GCL 68.55±0.68 81.56±1.31 68.66±3.13 75.79±1.97 54.68±0.82 58.33±4.46 54.20±4.83 55.30±3.48
SimGRACE 67.71±0.63 79.47±0.36 62.54±0.80 85.93±1.67 54.94±0.85 47.16±4.60 54.11±1.26 58.19±0.80
Gprompt DGI 61.28±6.92 72.35±6.43 60.00±4.15 59.82±8.05 53.15±3.14 49.01±1.79 50.69±8.24 52.10±4.06
GraphMAE 56.22±7.41 44.90±0.12 53.63±3.55 64.43±15.28 52.29±3.02 55.66±5.33 50.42±1.52 55.22±1.63
EdgePreGPPT 63.01±2.64 74.04±6.87 65.82±4.00 56.24±15.55 51.97±3.30 49.66±7.34 50.37±2.68 57.12±1.01
EdgePreGprompt 61.76±6.92 74.20±7.18 63.20±4.24 68.36±15.60 54.30±3.58 51.46±3.88 52.11±4.43 55.65±4.85
GCL 62.10±7.10 72.58±8.88 55.52±9.16 54.21±19.23 52.15±2.03 57.07±3.48 52.27±2.39 57.24±4.44
SimGRACE 63.49±7.53 73.85±5.74 58.07±7.27 56.06±18.58 53.72±4.10 53.97±4.68 49.80±7.28 56.78±6.19
GPF DGI 68.17±1.10 76.98±3.82 60.16±1.20 49.38±22.29 47.07±1.84 48.08±2.39 57.22±3.09 48.85±1.22
GraphMAE 68.64±0.80 50.00±0.00 60.27±1.58 75.23±2.86 52.79±1.96 52.96±5.14 59.08±2.59 57.07±5.22
EdgePreGPPT 72.39±1.43 78.68±3.14 59.70±1.16 71.74±1.36 53.23±0.59 55.95±3.53 57.16±1.15 55.12±1.37
EdgePreGprompt 64.81±4.48 76.99±3.95 59.65±1.05 75.64±1.21 55.95±1.04 61.22±3.30 57.14±3.83 60.99±3.42
GCL 68.50±0.89 79.36±0.64 58.20±2.58 75.42±2.40 55.17±1.18 57.66±3.49 56.26±0.72 54.38±2.29
SimGRACE 68.88±0.73 79.18±0.70 59.03±1.13 54.34±11.05 56.49±0.82 59.26±4.32 53.33±0.93 58.78±0.91
GPF-plus DGI 67.30±2.04 78.76±3.06 59.98±1.03 78.67±3.24 55.76±0.99 58.01±1.54 55.60±4.64 53.17±0.34
GraphMAE 66.86±1.34 50.00±0.00 61.20±1.57 72.94±0.86 49.96±0.80 47.18±2.64 54.12±12.01 59.32±3.11
EdgePreGPPT 68.32±1.40 79.72±3.57 61.31±2.41 68.83±3.06 55.99±0.45 61.66±1.49 51.94±1.06 49.50±2.07
EdgePreGprompt 66.08±4.10 78.31±2.86 58.73±2.17 51.73±17.01 57.28±2.04 59.88±2.24 52.69±2.74 55.53±2.48
GCL 67.76±1.67 78.53±0.70 57.90±2.40 77.86±2.73 53.24±1.85 61.30±2.87 50.97±1.94 52.80±2.36
SimGRACE 67.38±2.08 79.41±0.77 53.05±5.96 49.04±15.66 55.30±1.21 59.92±3.03 52.68±1.00 51.35±2.70
Table 26: Graph classification performance Accuracy (%) (5-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 62.60±4.01 55.23±4.26 62.90±5.03 73.47±3.92 25.67±0.48 64.99±10.42 51.48±2.29 63.59±2.86
Pretrain + Fine-tuning DGI 50.80±0.88 60.72±2.09 62.47±4.25 74.53±3.71 25.17±1.67 68.36±11.83 24.94±7.80 58.94±9.66
GraphMAE 55.75±5.48 54.40±5.15 61.21±6.92 67.33±5.61 26.13±1.09 68.42±11.71 38.27±18.11 64.71±3.22
EdgePreGPPT 60.97±10.07 57.44±3.75 62.85±4.63 70.80±3.05 27.04±1.75 69.38±10.91 58.40±12.79 64.52±2.97
EdgePreGprompt 62.88±4.02 58.11±3.76 60.25±4.96 73.20±2.44 27.46±1.29 73.19±9.53 34.20±16.57 62.70±3.37
GCL 65.40±3.33 48.28±6.30 63.33±4.13 75.33±1.89 24.63±1.52 55.44±3.84 72.96±11.98 63.84±2.06
SimGRACE 60.65±4.54 48.40±5.90 63.08±4.66 71.73±2.33 24.79±1.83 61.50±16.06 70.80±16.29 62.25±3.23
GPPTPrompt DGI 49.90±0.84 48.54±9.22 54.49±9.48 70.53±3.90 19.54±3.17 47.08±25.71 54.75±19.07 52.65±9.39
GraphMAE 49.60±0.75 33.61±16.35 58.27±4.63 69.20±3.30 22.17±2.34 56.14±25.93 69.63±14.96 57.07±5.16
EdgePreGPPT 66.37±3.59 50.16±9.23 57.37±4.85 65.33±3.65 21.33±1.04 60.05±13.16 45.37±19.12 51.36±4.26
EdgePreGprompt 49.60±0.75 37.57±10.07 46.92±11.86 40.67±14.47 21.96±2.12 67.88±17.34 48.02±12.75 60.02±3.24
GCL 59.58±9.19 55.21±1.21 58.25±3.13 70.40±4.10 20.92±2.40 64.50±19.35 56.11±21.68 57.05±3.15
SimGRACE 60.35±8.28 54.05±4.58 56.85±10.85 68.67±5.98 20.96±1.23 60.80±13.90 60.80±8.71 56.71±4.10
ALL-in-one DGI 60.47±7.38 57.59±5.73 71.37±4.89 84.00±2.11 26.46±2.24 48.15±4.32 61.73±8.94 54.59±2.82
GraphMAE 59.37±9.45 36.86±13.96 68.72±3.98 82.67±2.89 24.00±0.87 48.79±7.42 57.10±17.82 63.44±1.35
EdgePreGPPT 63.62±2.30 57.86±5.88 70.56±2.54 84.67±0.73 23.92±2.14 52.17±5.59 62.78±10.18 53.93±4.10
EdgePreGprompt 60.18±7.66 57.63±7.77 63.33±2.98 80.53±1.54 25.00±1.56 51.05±7.26 49.69±13.79 62.78±0.65
GCL 61.80±4.92 57.04±4.46 69.69±6.19 75.87±0.98 26.71±2.17 56.41±6.50 51.67±9.26 60.62±3.84
SimGRACE 63.05±3.01 54.01±0.83 68.72±4.97 75.47±3.08 26.67±0.99 62.95±8.57 53.64±5.97 59.32±0.34
Gprompt DGI 53.05±10.49 60.62±2.31 61.30±3.46 62.93±15.66 21.38±1.97 46.54±7.06 59.38±14.43 54.12±5.51
GraphMAE 55.15±7.10 32.55±0.12 58.00±7.16 65.87±8.50 19.87±2.24 52.23±5.40 53.77±6.74 50.68±5.33
EdgePreGPPT 66.70±3.87 57.23±2.43 62.94±1.38 63.87±12.45 17.79±1.27 53.35±4.42 51.98±8.17 52.46±4.48
EdgePreGprompt 52.75±10.26 57.71±1.56 55.55±6.18 73.07±2.13 20.79±2.66 52.12±7.21 54.32±11.18 56.24±3.69
GCL 61.38±10.05 60.71±4.54 60.54±2.22 64.40±8.95 20.33±2.27 53.24±10.24 56.73±7.80 58.28±2.18
SimGRACE 54.85±10.71 60.76±5.08 61.62±2.86 63.47±7.90 21.46±2.27 53.35±7.75 51.36±9.14 54.33±4.04
GPF DGI 59.90±9.95 59.65±6.25 60.99±4.00 72.93±3.14 24.75±0.98 52.98±15.39 61.05±11.51 58.70±1.84
GraphMAE 66.02±3.96 29.68±13.57 63.37±4.37 70.13±2.58 23.92±2.16 55.87±15.72 54.63±9.43 60.57±3.74
EdgePreGPPT 67.80±5.58 58.07±6.03 63.28±4.33 68.40±3.44 24.13±1.56 66.27±14.57 32.53±10.35 55.73±3.64
EdgePreGprompt 62.62±4.73 56.62±5.44 62.34±3.36 74.00±3.65 24.29±1.46 64.77±7.22 50.99±16.41 61.06±2.63
GCL 60.17±9.46 56.65±2.44 62.34±4.37 70.67±3.89 27.00±0.78 58.28±9.07 53.70±10.84 59.07±0.65
SimGRACE 62.33±4.53 55.67±7.42 60.74±3.46 72.27±4.51 26.21±0.96 29.81±15.53 27.10±4.50 43.52±5.74
GPF-plus DGI 63.50±5.68 54.91±7.08 63.51±2.89 71.87±5.55 25.67±1.75 60.16±3.91 62.84±9.18 62.76±4.22
GraphMAE 63.72±5.43 37.23±18.02 63.15±4.75 72.67±1.89 24.79±1.80 67.35±13.04 49.51±9.62 59.21±6.53
EdgePreGPPT 68.13±3.31 60.68±4.67 61.89±4.59 70.93±3.00 23.04±1.13 68.20±12.65 24.01±5.97 64.46±3.57
EdgePreGprompt 63.85±5.88 57.76±6.85 58.83±2.28 72.67±4.04 26.87±1.89 72.87±10.17 22.35±2.72 64.80±3.45
GCL 60.18±10.28 57.43±3.27 63.08±4.23 73.87±3.51 25.79±1.76 60.70±11.73 57.59±5.89 52.78±10.34
SimGRACE 60.43±9.87 58.99±4.32 63.10±4.06 72.80±3.30 24.58±1.68 61.77±16.67 71.54±14.81 61.97±3.07
Table 27: Graph classification performance F1-score (5-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 61.53±3.82 46.72±7.81 60.12±2.97 68.77±4.51 21.36±3.02 50.40±4.10 49.85±2.02 61.16±2.22
Pretrain + Fine-tuning DGI 37.27±7.49 60.77±1.97 59.48±2.21 69.60±5.01 23.93±3.09 46.29±3.39 21.92±9.07 53.50±13.04
GraphMAE 48.52±12.68 47.93±10.31 58.76±5.50 61.19±5.53 21.10±2.81 46.30±3.29 34.64±16.14 60.62±2.52
EdgePreGPPT 59.95±10.96 55.01±7.38 58.97±1.92 66.40±3.88 23.13±4.27 46.90±4.51 49.13±5.88 60.32±3.32
EdgePreGprompt 61.87±3.87 58.08±3.74 57.72±3.83 67.43±5.46 24.23±3.11 45.22±2.84 29.92±15.69 55.91±9.68
GCL 64.48±3.90 46.25±8.10 60.16±1.94 72.18±2.14 20.23±2.20 50.69±2.30 44.56±0.94 61.72±1.92
SimGRACE 58.80±4.71 46.94±6.45 60.31±2.61 67.55±3.52 19.51±3.69 50.79±10.90 42.64±2.95 54.10±1.53
GPPTPrompt DGI 33.29±0.37 44.26±14.53 51.17±11.32 62.76±9.57 15.91±3.76 31.80±11.62 40.57±12.66 49.10±8.69
GraphMAE 33.15±0.33 16.01±6.20 56.63±5.43 61.64±4.83 18.70±2.65 36.00±12.12 45.03±4.50 48.97±2.70
EdgePreGPPT 65.92±3.82 44.90±15.76 39.47±4.13 60.33±5.13 17.40±1.52 45.39±5.02 38.16±10.87 50.83±4.76
EdgePreGprompt 33.15±0.33 19.75±6.76 44.10±11.48 33.67±17.23 19.46±2.35 42.06±4.74 41.95±7.64 55.63±2.35
GCL 55.16±14.72 55.24±1.26 40.33±5.84 65.16±6.09 19.05±1.73 46.02±10.48 39.32±8.94 54.10±2.94
SimGRACE 59.66±9.14 54.05±3.81 54.48±10.12 62.48±5.78 19.62±1.57 49.41±6.89 50.14±3.71 54.01±3.13
ALL-in-one DGI 59.53±8.82 57.32±5.81 70.60±4.58 82.31±2.87 17.66±2.59 44.71±2.85 52.91±7.30 52.63±1.53
GraphMAE 58.72±10.43 17.43±5.21 66.84±3.16 80.56±3.23 15.28±0.64 45.02±5.39 50.25±14.39 57.48±1.16
EdgePreGPPT 63.43±2.34 57.44±6.03 69.85±2.07 83.59±1.08 18.20±5.07 47.22±3.60 46.61±3.47 53.27±3.89
EdgePreGprompt 58.91±9.07 56.52±8.78 61.72±1.75 78.93±1.51 19.95±1.44 48.65±5.08 44.03±8.76 59.90±1.32
GCL 61.47±5.33 56.68±4.02 68.30±5.95 73.50±0.83 22.30±0.99 47.14±2.57 45.65±4.16 53.88±4.86
SimGRACE 62.85±2.99 55.22±0.57 67.22±4.47 73.69±2.49 21.06±0.77 51.71±4.63 48.74±3.72 40.65±6.82
Gprompt DGI 50.47±12.00 60.36±1.92 58.82±4.98 59.86±14.59 19.68±2.10 43.07±4.95 52.76±11.53 52.64±4.28
GraphMAE 51.26±10.62 16.37±0.04 56.08±6.65 61.49±7.64 18.48±2.23 46.44±3.42 48.92±4.67 49.81±4.74
EdgePreGPPT 66.49±4.14 57.50±2.41 61.89±1.74 61.80±11.24 16.31±1.68 48.08±4.53 46.79±5.22 51.83±4.02
EdgePreGprompt 50.18±11.72 58.02±1.57 53.81±6.27 70.82±1.39 18.42±3.25 46.63±4.27 47.81±6.75 55.47±3.78
GCL 60.64±11.30 59.83±4.00 57.86±1.00 60.49±10.88 19.29±2.26 47.44±7.08 50.30±4.88 57.01±2.25
SimGRACE 53.16±12.04 60.09±4.57 59.49±2.71 59.62±6.15 20.75±2.19 49.12±4.71 47.73±7.10 52.97±4.17
GPF DGI 58.67±10.64 59.18±6.42 58.89±2.28 67.69±4.30 18.41±2.04 43.64±5.83 56.38±8.64 43.94±8.16
GraphMAE 65.59±4.08 14.71±5.24 60.67±2.38 65.88±1.64 19.47±2.56 45.45±8.77 49.37±6.61 57.04±2.26
EdgePreGPPT 67.73±5.58 57.75±6.17 61.14±4.21 63.56±4.73 19.89±2.46 43.97±2.80 30.06±8.48 55.24±3.51
EdgePreGprompt 61.77±4.51 56.12±5.23 60.54±2.00 72.05±3.02 19.31±2.59 50.99±3.91 39.07±4.52 58.14±3.07
GCL 58.58±10.31 54.42±4.20 58.69±2.31 67.61±3.98 21.36±3.31 50.58±4.57 47.55±4.52 40.86±7.24
SimGRACE 61.27±4.49 54.22±8.92 59.44±2.34 67.56±6.57 23.55±3.05 25.14±14.15 26.14±5.64 33.65±9.49
GPF-plus DGI 63.03±5.96 53.58±8.50 59.69±2.41 70.16±4.83 21.63±0.48 52.70±2.44 57.31±6.19 57.95±3.57
GraphMAE 62.85±5.85 17.18±6.87 60.52±2.74 67.74±2.44 21.75±1.83 45.30±2.36 45.36±7.65 56.39±7.29
EdgePreGPPT 67.66±3.65 60.45±4.30 58.78±2.63 68.26±2.34 20.46±1.22 45.80±4.06 21.07±7.39 60.10±2.62
EdgePreGprompt 63.38±6.30 56.87±7.08 55.28±2.90 69.63±3.06 23.58±3.05 44.79±1.97 19.35±3.96 60.64±2.13
GCL 58.81±10.82 57.11±3.43 60.02±2.04 69.79±2.97 22.03±2.03 47.74±4.03 53.04±4.37 47.52±13.56
SimGRACE 59.10±10.57 58.56±4.61 60.30±2.04 69.16±2.98 22.60±2.36 51.12±11.99 43.15±1.95 53.98±1.48
Table 28: Graph classification performance AUROC (5-shot)
Training schemes Methods IMDB-B COLLAB PROTEINS MUTAG ENZYMES COX2 BZR DD
self-supervise GCN 66.66±4.29 65.98±11.45 60.55±1.27 68.80±5.31 59.06±0.90 53.16±1.88 63.16±1.91 64.75±3.62
Pretrain + Fine-tuning DGI 50.02±0.04 81.66±0.61 60.91±1.77 68.55±5.83 58.93±1.27 51.76±2.64 52.81±2.18 63.61±4.41
GraphMAE 57.68±8.77 69.01±10.12 59.92±2.11 74.33±4.22 58.39±0.82 51.25±2.21 57.13±7.05 64.45±3.57
EdgePreGPPT 61.63±14.66 76.71±9.50 60.88±1.39 72.84±3.58 59.64±0.79 53.48±3.98 60.58±6.74 63.82±3.08
EdgePreGprompt 66.78±4.49 81.42±0.65 56.93±4.04 72.62±1.98 59.65±1.08 51.34±2.69 50.75±1.31 62.82±3.74
GCL 69.68±5.15 76.71±2.70 61.80±2.09 70.63±4.24 57.49±0.69 55.91±4.67 50.86±1.72 64.37±3.44
SimGRACE 64.69±7.42 72.81±3.58 61.49±2.04 71.48±3.93 55.98±2.05 61.95±2.47 48.92±2.16 57.39±2.98
GPPTPrompt DGI 50.48±0.70 70.62±7.22 55.92±14.32 79.97±4.66 52.20±2.26 51.22±2.96 47.75±9.82 53.81±9.82
GraphMAE 50.63±0.56 50.56±3.78 60.56±6.96 77.76±5.00 53.63±1.06 50.60±6.83 53.55±6.08 58.60±5.56
EdgePreGPPT 69.02±4.94 68.78±9.45 68.50±9.36 72.44±6.57 53.43±1.03 53.02±8.32 47.93±5.96 51.76±5.53
EdgePreGprompt 50.63±0.56 56.85±1.07 45.95±18.22 24.75±25.84 53.66±1.73 52.56±3.29 49.31±5.04 61.01±3.67
GCL 59.75±10.88 76.73±1.15 69.17±8.03 77.80±5.27 53.37±1.20 56.22±8.12 50.45±6.37 57.89±3.91
SimGRACE 61.35±10.03 70.93±4.01 59.76±13.58 77.27±7.67 53.63±1.06 57.78±6.95 59.49±6.13 57.49±4.92
ALL-in-one DGI 61.66±12.01 80.99±1.96 79.39±1.76 90.73±0.46 58.71±0.88 49.33±1.56 52.97±11.15 52.92±1.58
GraphMAE 60.44±13.36 50.00±0.00 73.82±1.89 89.59±1.17 53.75±1.13 54.24±10.01 55.73±11.76 64.55±1.63
EdgePreGPPT 69.53±3.98 81.50±1.45 77.60±0.81 91.73±0.49 54.18±3.09 52.22±5.14 39.69±2.87 58.41±0.60
EdgePreGprompt 61.78±11.31 81.93±1.28 66.26±1.84 87.58±0.76 55.09±2.15 61.86±2.18 50.64±4.57 64.95±0.92
GCL 65.55±5.57 77.86±4.19 73.65±4.67 85.46±1.40 59.50±1.10 49.77±1.51 52.63±2.91 59.95±4.37
SimGRACE 66.52±3.18 75.84±0.40 77.13±0.44 84.04±2.22 55.76±0.82 62.42±1.89 54.60±2.87 55.43±0.66
Gprompt DGI 53.91±13.03 70.52±5.99 57.55±7.36 64.03±19.95 53.87±2.02 50.06±2.93 56.36±6.92 54.21±4.23
GraphMAE 55.46±7.98 44.83±0.15 54.75±7.17 63.37±9.48 52.82±1.11 49.40±6.97 54.99±5.88 51.72±5.54
EdgePreGPPT 70.26±4.44 73.55±4.01 67.07±0.80 66.04±10.33 53.32±2.77 52.47±6.25 52.16±2.83 53.26±4.39
EdgePreGprompt 56.66±8.48 75.59±3.42 56.24±7.10 74.43±4.69 54.57±2.69 50.60±3.01 53.90±6.28 56.36±5.42
GCL 52.62±12.32 75.45±3.52 61.03±5.77 64.43±10.15 54.04±2.81 50.64±7.63 53.71±2.16 56.22±3.97
SimGRACE 57.24±12.21 73.41±3.61 60.58±5.34 64.67±8.28 55.02±1.57 56.79±5.66 56.83±6.49 53.99±6.27
GPF DGI 61.28±14.33 80.91±1.35 59.36±1.27 68.29±4.20 57.38±0.69 55.65±2.89 69.56±8.43 52.10±1.18
GraphMAE 71.36±4.86 50.00±0.00 60.79±1.35 73.75±6.18 57.51±1.83 55.80±3.40 57.87±6.08 59.78±6.01
EdgePreGPPT 72.09±6.35 81.60±1.83 63.33±6.33 71.14±4.08 57.09±1.47 45.44±3.51 40.30±2.41 57.56±1.55
EdgePreGprompt 66.56±4.93 82.11±0.58 59.32±1.46 78.72±2.44 56.88±0.88 57.11±3.21 38.86±3.08 59.75±5.73
GCL 61.77±14.84 76.32±2.61 58.70±2.93 75.54±4.82 57.23±1.32 50.87±1.24 54.74±5.41 51.02±2.03
SimGRACE 66.69±4.51 78.63±3.50 60.84±0.83 73.60±4.98 57.64±0.64 55.74±2.01 44.52±2.78 48.14±1.17
GPF-plus DGI 67.76±7.65 80.06±1.94 60.56±1.11 75.93±5.59 57.99±1.37 56.06±2.35 67.92±4.56 62.75±3.59
GraphMAE 68.56±7.52 50.00±0.00 60.88±1.41 74.54±6.62 56.78±1.51 48.93±2.41 54.11±8.51 62.56±3.54
EdgePreGPPT 71.71±3.89 82.51±0.50 59.30±1.96 77.03±2.00 56.17±0.68 49.53±4.57 49.78±0.43 63.49±4.00
EdgePreGprompt 67.36±7.32 81.68±0.93 54.63±1.53 75.31±2.84 60.04±0.95 50.95±1.90 48.60±2.80 64.20±3.47
GCL 61.80±14.82 80.57±0.98 61.23±1.64 69.47±3.56 57.95±1.27 51.87±1.80 64.97±3.46 61.94±4.01
SimGRACE 61.85±14.76 77.71±3.28 61.07±1.57 70.13±2.87 57.30±0.70 63.15±2.98 49.11±1.78 57.21±2.94
Figure 5: The curve of the shot number from 1 to 10 and the accuracy (%) on Cora and ENZYMES.
Table 29: Node classification performance Accuracy (%) across different graph models on Wisconsin (1-shot setting). The Accuracy of supervised methods: GAT is 34.51±18.02, GraphSAGE is 25.37±5.61, GT is 20.91±7.07
Pre-train & Fine-tune
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 16.00±6.24 37.60±10.69 20.00±3.82 33.37±4.76 18.86±1.88 28.00±9.40
GraphSAGE 40.69±9.46 43.77±12.43 26.06±5.38 29.94±3.75 36.57±4.88 9.37±2.72
GT 25.71±3.07 39.77±8.42 23.20±2.65 28.23±6.64 11.77±1.06 14.51±5.08
GPPT
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 22.17±6.13 33.94±7.76 23.43±4.46 37.94±7.11 26.86±6.12 29.83±8.04
GraphSAGE 26.51±8.00 30.51±5.40 21.49±5.17 24.23±6.55 20.91±7.11 25.37±7.22
GT 27.20±10.48 29.83±5.80 28.00±6.01 23.31±3.01 27.66±0.69 25.03±5.43
All-in-one
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 69.44±5.19 36.25±10.63 91.25±4.33 92.65±3.75 42.85±9.16 36.61±14.86
GraphSAGE 94.73±5.86 97.55±0.78 98.60±0.87 100.00±0.00 83.01±18.37 85.12±16.29
GT 68.59±11.78 98.48±0.70 96.74±1.01 100.00±0.00 70.23±8.80 45.70±11.83
Gprompt
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 58.25±13.83 67.77±15.91 94.17±2.26 84.28±3.63 80.11±16.65 57.18±12.60
GraphSAGE 66.48±12.88 83.49±15.93 87.52±3.79 82.16±2.64 65.50±6.48 72.61±5.97
GT 56.03±7.33 73.50±9.72 76.97±13.39 80.07±2.84 59.31±10.17 69.30±10.57
GPF
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 63.64±3.91 78.59±18.57 97.90±0.79 91.84±5.19 74.93±4.49 65.67±15.07
GraphSAGE 68.12±13.96 67.66±13.37 74.06±14.59 72.45±10.14 59.69±21.37 78.37±14.84
GT 39.85±4.83 71.26±14.43 72.67±13.36 81.33±3.41 78.19±2.19 67.90±10.53
GPF-plus
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 92.62±8.60 96.37±4.44 99.29±1.41 97.88±4.24 86.12±7.73 94.17±2.63
GraphSAGE 77.85±3.30 87.88±0.90 99.30±0.86 96.16±5.09 85.93±19.52 72.47±8.33
GT 91.69±6.84 98.60±1.02 99.30±0.86 96.13±4.27 58.83±18.53 89.04±4.25
Table 30: Graph classification performance Accuracy (%) across different graph models on PROTEINS (1-shot setting). The Accuracy of supervised methods: GATis 48.34±9.96, GraphSAGE is 60.54±2.95, GT is 61.44±2.48
Pre-train & Fine-tune
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 58.34±6.52 61.06±4.13 63.75±3.71 54.09±4.03 60.04±3.06 58.65±6.71
GraphSAGE 60.70±4.08 60.56±5.12 61.60±1.78 63.21±1.80 61.80±3.77 58.56±1.84
GT 53.87±4.81 60.00±3.99 64.92±3.19 56.58±3.28 62.88±1.82 60.00±1.60
GPPT
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 57.71±8.98 57.80±10.55 58.04±9.92 54.97±7.45 52.29±7.83 55.15±9.84
GraphSAGE 56.56±6.73 57.73±7.95 58.63±11.78 56.94±5.67 58.00±7.80 54.74±6.59
GT 53.08±7.56 57.35±8.58 60.27±3.92 55.51±7.68 56.18±5.79 55.87±7.69
All-in-one
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 60.04±3.84 60.00±6.04 62.11±2.85 63.21±2.22 58.36±4.93 59.37±5.59
GraphSAGE 59.53±4.94 60.70±4.89 63.12±1.59 59.98±8.46 62.22±3.81 62.04±2.07
GT 57.39±3.66 58.92±6.61 62.61±4.08 60.20±7.55 62.81±1.63 50.52±6.17
Gprompt
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 61.08±6.19 63.03±2.61 64.47±4.30 61.48±3.34 59.12±6.84 58.13±7.27
GraphSAGE 61.35±2.21 59.48±9.19 60.92±3.16 63.30±1.43 55.26±2.61 63.21±2.66
GT 56.65±5.81 60.99±1.62 61.87±5.60 55.33±3.69 54.81±7.62 58.97±1.16
GPF
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 66.11±5.18 65.62±5.38 56.38±8.64 65.55±6.65 59.71±7.66 67.42±6.26
GraphSAGE 64.20±7.63 67.87±4.32 58.18±9.06 64.49±6.80 60.25±2.91 62.94±2.29
GT 65.80±7.42 60.16±5.81 64.54±7.18 61.21±2.91 58.74±5.51 59.57±2.93
GPF-plus
Model DGI GraphMAE EdgePreGPPT EdgePreGprompt GCL SimGRACE
GAT 56.20±12.87 57.35±11.28 56.25±8.61 53.24±4.79 57.48±11.74 57.48±9.63
GraphSAGE 56.22±9.08 57.55±10.56 56.31±9.26 57.71±9.60 53.89±9.47 55.89±4.30
GT 53.39±5.23 57.37±10.95 57.39±11.88 52.61±5.30 57.62±12.27 56.16±5.07
(a) GPPT(Node Task)
(b) All-in-one(Node Task)
(c) Gprompt(Node Task)
(d) GPF(Node Task)
(e) GPPT(Graph Task)
(f) Gprompt(Graph Task)
(g) GPF(Graph Task)
(h) GPF-plus(Graph Task)
Figure 6: Heat maps of graph prompt methods on different node/graph-level datasets in 1-shot settings, excluding GPF-plus on the node task and All-in-one on the graph tasks.